### 24 Jan 2024 v1

# Functoriality in Finitary Vacuum Einstein Gravity and Free Yang-Mills

Abstract
We continue ongoing research work [53, 54, 55, 67, 68, 69, 70, 71, 72, 73, 74, 75] on applying the homological algebraic conceptual and technical machinery of Abstract Differential Geometry (ADG) towards formulating a finitary, causal and quantal version of vacuum Einstein Lorentzian gravity and free Yang-Mills theories, hitherto cumulatively referred to as ADG-Gauge Theory (ADG-GT). In particular, we unfold, express and highlight the inherently functorial character of ADG-GT both at the ‘kinematical’ and at the ‘dynamical’ level of the aufbau of the theory, although at the same time we observe that the traditional kinematics-versus-dynamics distinction becomes blurry in our ADG-theoretic approach as, in line with [55, 56, 70, 72, 73, 61], we maintain that there is no pre-existent geometrical/kinematical space in the quantum deep, but rather, that physical geometrical space derives from (or is an outcome of) field dynamics. We moreover argue that the gauge theory of the third kind and the third quantisation schemes that ADG-GT has been seen to support [55, 72, 73], are also functorial in character. Furthermore, since our inherently algebraic ADG-theoretic scheme has been seen to be manifestly background geometrical $\mathcal{C}^{\infty}$-smooth spacetime manifold independent [53, 54, 55, 67, 68, 69, 70, 71, 72, 73], we entertain the idea, at the ‘dynamical’ level of functoriality, that there is both a geometric morphism and a natural transformation type of correspondences between the relevant Einstein and Yang-Mills field functor categories with the dynamical gauge connection and curvature sheaf morphisms implementing the homological algebraic dynamics within each, so as to further corroborate previous claims [69, 70, 71, 72, 73] that from an ADG-theoretic perspective, ADG-gravity is an already finitistic, third quantised, $\mathcal{C}^{\infty}$-smooth geometrical background spacetime manifoldless, auto-dynamical and ‘pure gauge’ field theory of the third kind. We also cast our formal canonical sheaf cohomological Third Quantisation heuristics originally formulated in [72] in a slightly different light so as to arrive at a new ADG-theoretic notion of ‘Unitary’ Quantal ADG-Gauge Field which, in a tetrad of functorially and dynamically entwined structures $\mathbf{U}:=(\mathbf{\mathcal{E}},\mathcal{D},{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}},\mathcal{Q})$, it subsumes under a single coherent and inseparable ‘unitary whole’ all the four most important functorial structural traits of ADG-GT: ‘local quantum particle states’ represented by local sections of a vector sheaf $\mathbf{\mathcal{E}}$, their ‘dual-complementary’ functorial ADG-gauge field dynamics generated by an algebraic $\mathbf{A}$-connection $\mathcal{D}$, the latter’s local gauge invariance of the 3rd kind encoded in the principal structure sheaf ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ of $\mathbf{\mathcal{E}}$’s automorphisms, and the dual particle-field canonical-type of 3rd quantisation, represented by the functorial morphism $\mathcal{Q}$ between the sheaf categories involved. At the end, we give a subjective (:from this author’s viewpoint) account of certain key ideas, concepts and seminal mathematical results in the past that significantly motivated Professor Mallios to develop ADG, and how these ideas subsequently inspired this author to apply it to a finitistic and quantal theoretical scenario for Vacuum Einstein Gravity and Free Yang-Mills theories. Throughout the second half of the paper, we recall and analyse several pertinent quotes that Professor Mallios and this author used to repeatedly discuss and scrutinise in the course of the late nineties and early noughties, during endless late night discussions over good food and wine at our favourite tavern, fittingly called Algebra, in Paleo Psychiko, Athens, Greece. An Addendum at the end of the paper recalls an important and telling early interaction that this author enjoyed with Professor Mallios at the end of last century.
PACS numbers: 04.60.-m, 04.20.Gz, 04.20.-q
Key words: functoriality, quantum gravity, quantum Yang-Mills theories, causal sets, differential incidence Rota algebras of locally finite partially ordered sets, finitary spacetime sheaves, abstract differential geometry, sheaf theory, sheaf cohomology, category theory, topos theory, geometric prequantisation, canonical quantisation
1 Technical Prolegomena: A Brief History of Finitary ADG-Gravity and Yang-Mills Theories with an Emphasis on the Functorial Character of our Concepts, Methods and Constructions
In this section we give a brief account of the main milestones reached along our way towards arriving at a purely algebraic, finitistic, causal and quantal theory of spacetime, gauge theories and gravity. As we outline the main results, we highlight and emphasise the homological algebraic (:category-theoretic), and especially functorial, nature of our basic concepts, structures and methods of their use in various constructions and associated (abstract differential geometric) calculations (:Differential Calculus), while all this is accomplished purely algebraically, manifestly without any recourse to or dependence on a background $\mathcal{C}^{\infty}$-smooth geometrical base spacetime manifold.
1.1 Finitary Substitutes of Continuous Spacetime Manifolds, their Incidence Algebras, and the Finitary Sheaves Thereof: ‘Kinematical’ Functoriality
1.1.1 Sorkin’s Finitary Posets
Our journey begins with Sorkin’s ‘prophetic’ finitary substitutes of continuous spacetime manifolds [78] [^1] Refer to this paper for various mathematical concepts, structures and technical definitions thereof.. In that paper, with every locally finite (:finitary) open cover $\mathcal{U}_{i}=(U_{i})$ of a (real) topological ($\mathcal{C}^{0}$) spacetime manifold $M$, Sorkin assigns a so-called finitary partially ordered set (finposet) $\mathcal{P}_{i}$:
$\mathcal{U}_{i}\longrightarrow\mathcal{P}_{i}\;\;\;$ (1)
The collection $\overleftarrow{\mathcal{P}}=(\mathcal{P}_{i})_{i\in I}$ of such finposets is seen to constitute a so-called inverse or projective system, or net, of posets, which is seen to have a projective limit space effectively homeomorphic to the continuous $\mathcal{C}^{0}$-manifold $M$ [^2] The index $i\in I$ in $\mathcal{P}_{i}$ is the so-called refinement net index, whereby $\mathcal{U}_{i}\prec\mathcal{U}{j}$ (reads: the open cover $\mathcal{U}{i}$ is finer than the open cover $\mathcal{U}{j}$, and conversely, $\mathcal{U}{j}$ is coarser than $\mathcal{U}{i}$) if $\mathcal{U}_{i}$ has ‘smaller’ and more numerous open sets than $\mathcal{U}_{j}$. The aforementioned inverse limit now reads: the $\mathcal{C}^{0}$-manifold is recovered at the inverse limit of infinite refinement of the open covers $\mathcal{U}_{i}$s in the net (as $i\rightarrow\infty$).
Three bullet points must be emphasised here in connection with (1) from [78]:
  • First, Sorkin’s original intuition that every geometrical point of a point-manifold is an ideal, operationally unrealistic and physically untenable plus problematic (:singular), dimensionless and structureless object, that better be ‘smeared’ and blown-up by ‘enlarged’ open sets (neighbourhoods) about it. [^3] Let it be noted here that a pointed background geometrical spacetime continuum is a problematic, pathological structure that is arguably responsible both for the singularities plaguing General Relativity (GR) and for the pestilential non-renormalisable unphysical infinities marring Quantum (Gauge) Field Theories (QFT) of matter. The open sets and their set-theoretic algebra are the carriers of the manifold’s topology and its continuity, not its ideal points. [^4] In the ensuing discussion, by introducing sheaf theory, we will refine this statement even more.
  • Sorkin’s original idea strongly resonates with Grothendieck’s pioneering idea to categorically abstract and generalise pointed topological spaces to pointless ones called sites by abstracting from the usual topological open covers like in $\overleftarrow{\mathcal{P}}=(\mathcal{U}_{i})=(\mathcal{P}_{i})_{i\in I}$ to families (=sieves) of covering arrows in a category defining a Grothendieck topology on the category [35, 36, 37]. [^5] A site is by definition a category endowed with a Grothendieck topology.
  • Once we have done away with the pointed geometrical manifold continuum, we can work further with their poset substitutes $(\mathcal{P}_{i})$ and build on them. That is what we do next.
First encounter with ‘Kinematical’ Functoriality. Before we go on to work with the finitary posets, we catch a first glimpse of functoriality of our constructions. Let $\mathcal{T}_{i}=\mathrm{span}\{U:\;U\in\mathcal{U}_{i}\}$ be the topology ‘spanned’ or generated by arbitrary unions and finite intersections of the open sets in each locally finite open covering $\mathcal{U}_{i}$. Then, a continuous map
$f:\;\mathcal{T}_{i}\longrightarrow\mathcal{T}_{j}$ (2)
induces (:maps functorially to) a poset morphism $\hat{f}:\;\mathcal{P}_{i}\rightarrow\mathcal{P}_{j}$, [^6] By definition, a poset morphism is a partial order preserving map. such that the following diagram commutes:
${\mathcal{T}_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${\mathcal{T}_{j}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{P}_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\hat{f}}$${\mathcal{P}{j}}$ (3)
1.1.2 Differential Incidence (Rota) Algebras of Finitary Posets and their Simplicial Complexes
From [94, 95, 74, 75] we read that with every finitary poset $\mathcal{P}{i}$ one can straightforwardly associate a finitary simplicial complex $\mathcal{S}_{i}$, the so-called Čech-Alexandrov nerve of the underlying finitary open covering, á-la Čech Homology:
$\mathcal{S}_{i}:\>\mathcal{P}_{i}\longrightarrow\mathcal{S}_{i}$ (4)
Second encounter with ‘Kinematical’ Functoriality. The mapping $\mathcal{S}$ above is also functorial in the sense that a finitary poset morphism $p_{ij}:\;\mathcal{P}_{i}\rightarrow\mathcal{P}_{j}$ as before, induces (:maps functorially to) a simplicial mapping: $\hat{s}_{ij}:\;\mathcal{S}_{i}\rightarrow\mathcal{S}_{j}$, so that the following diagram commutes:
${\mathcal{P}_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{S}_{i}}$$\scriptstyle{p_{ij}}$${\mathcal{P}_{j}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{S}_{j}}$${\mathcal{S}_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\hat{s}_{ij}}$${\mathcal{S}_{j}}$ (5)
Third encounter with ‘Kinematical’ Functoriality: Incidence (Rota) Algebras. More importantly for our considerations here, we read from [94, 95, 74, 75] that with every finitary poset $\mathcal{P}_{i}$ (or equivalently, with every finitary Čech simplicial complex $\mathcal{S}_{i}$) one can naturally associate a so-called incidence (Rota) algebra $\Omega_{i}$ over the complex numbers $\mathbb{C}$, as follows:
$\mathcal{R}_{i}:\;\mathcal{P}_{i}\rightarrow\Omega_{i}$ (6)
Like in the case of the simplicial mapping of finitary posets $\mathcal{S}$, the mapping $\mathcal{R}$ induces an incidence algebra homomorphism $\hat{r}$ in the following functorial commutative diagram sense: [^7] Mutatis mutandis for the finitary Čech simplicial complexes and the incidence Rota algebras thereof: their correspondence is manifestly functorial [94, 95, 74, 75].
${\mathcal{P}_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{R}_{i}}$$\scriptstyle{p_{ij}}$${\mathcal{P}_{j}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{R}_{j}}$${\Omega_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\hat{r}_{ij}}$${\Omega_{j}}$ (7)
1.1.3 Fourth encounter with ‘Kinematical’ Functoriality: Gel’fand Duality
Without going into detailed technicalities here, we read from [94, 95, 74, 75] that one can go the other way around and extract from the finitary incidence algebras a topological space, endowed with a so-called Rota topology, by considering irreducible representations of the incidence algebras, the kernels of which correspond to primitive ideals in the algebras. In turn, the set of primitive ideals, the so-called spectrum of the algebra $\mathrm{Spec}(\Omega)$, now regarded as a generalised, ‘blown up point set’, is readily endowed with a Rota topology in such a way that the incidence algebra homomorphisms in ${\hat{r}_{ij}}$ lift to continuous maps in the respective Rota topologies.
This is another instance of the functoriality of our constructions and it corresponds to a finitary version of Gel’fand Duality according to which, very broadly speaking, from an algebraic structure $A$ one can extract a ‘geometrical space’ $\mathrm{Spec}(A)$ carrying a ‘natural continuity’ (:a functorially imposed topology on it). [^8] Furthermore, and again very broadly speaking, there is the Gel’fand Representation theorem that ensures that the $\mathbb{C}$-algebra of continuous complex valued functions on $\mathrm{Spec}(A)$ is naturally equivalent to the (complex) algebra $A$ that one started with. This remark will prove crucial in the sequel when we recount the introduction of finitary spacetime sheaves.
1.1.4 Finitary Incidence (Rota) Algebras as Finitary Differential Algebras/Modules
We read again directly from [74, 75, 54] that the finitary incidence algebras $\Omega_{i}$ are $\mathbb{Z}_{+}$-graded discrete differential algebras/modules of finite rank, [^9] See also [10, 11, 12] for an early study of such differential spaces. as follows:
$\Omega=\bigoplus_{n\in\mathbb{Z}_{+}}\Omega^{n}=\Omega^{0}\oplus\Omega^{1}\oplus\ldots=\mathbf{A}\oplus\mathcal{R}$ (8)
The $\Omega^{n}$s above are seen to be the reticular analogues of the usual linear spaces of $n$-grade (Grassmann exterior) differential forms [54] on a $\mathcal{C}^{\infty}$-smooth manifold. The grade 00 commutative linear subalgebra $\mathbf{A}=\Omega^{0}$ is the discrete analogue of the algebra of (smooth) functions (00-forms) on the continuum, while $\mathcal{R}=\bigoplus_{n\geq 1}\Omega^{n}$ serves as the $\mathbf{A}$-module of discrete differential forms on it.
Furthermore, we witness in [54, 55] that there is a discrete analogue of the (flat) Cartan-Kähler (exterior) differential $d$ operator:
$d:\;\Omega^{n}\longrightarrow\Omega^{n+1}$ (9)
that is a linear map and it obeys the Leibniz rule.
1.1.5 The Differential Caveat: Finitary Spacetime Sheaves of Incidence Algebras and Preliminary Vibes of ADG
As soon as this author realised that the incidence algebras encode not only topological, but also differential geometric, information in their structure, the next tenable position would be to somehow make them (dynamically) variable, thus he envisaged to employ in the longer run the full ADG-theoretic panoply towards formulating a finitary and quantal version of Gravity and Gauge Theory.
A first step to that end would be to ‘sheafify’ them; that is, to consider sheaves thereof. [^10] Along very similar lines of thought, the reader should refer to the Introduction of [61] to read how the notion of a sheaf comes hand in hand with the notion of variable structure. Thus finitary spacetime sheaves [68] $\mathbf{\mathcal{E}}_{i}$ of incidence algebras [^11] Now the locally finite posets being interpreted as causal and quantal versions of Sorkin’s causal sets [5, 79, 80, 77], as expounded in [67]. over Sorkin’s finitary poset discretisations were born.
Fifth encounter with ‘Kinematical’ Functoriality: Sheafification. That is, this author realised in [68] that the mapping $\mathcal{R}_{i}:\;\mathcal{P}_{i}\rightarrow\Omega_{i}$ in 6 above is actually a contravariant functor [^12] Effectively, the definition of a presheaf [6]. which, when subjected to suitable compatibility (glueing) conditions, it can be promoted to a local homeomorphism (between the corresponding covering topologies $\mathcal{T}_{i}$ generated by the finitary open covers $\mathcal{U}_{i}$), the very definition of a sheaf [6].
All in all, when the resulting finitary spacetime sheaves $\mathbf{\mathcal{E}}_{i}$ have $\Omega_{i}$s in their stalks, they were recognised as being the reticular analogues of Mallios’s vector sheaves. [^13] By definition, Mallios’s vector sheaves are locally free (differential) $\mathbf{A}$-modules of finite rank $n$ [38, 39, 44]. That is, locally for every open set $U\in\mathcal{U}{i}$ in an open covering (:set of local gauges $\mathcal{U}_{i}$) of the base topological space $X$ of a $\mathbb{C}$-algebraized space $(X,\mathbf{A})$, one has by definition the following $\mathbf{A}|_{U}$-isomorphisms: $\mathbf{\mathcal{E}}|_{U}=\mathbf{A}^{n}|_{U}=(\mathbf{A}|_{U})^{n}$ and, concomitantly, the following equalities section-wise: $\mathbf{\mathcal{E}}(U)=\mathbf{A}^{n}(U)=\mathbf{A}(U)^{n}$ (with $\mathbf{A}^{n}$ the $n$-fold Whitney sum of $\mathbf{A}$ with itself). Hence, the whole enterprise of applying Mallios’s Abstract Differential Geometry (ADG) to to a finitistic, causal and quantal version of Lorentzian vacuum Einstein Gravity and free Yang-Mills (gauge) theories of matter commenced.
In this line of thought, in [53, 54, 55] we defined finitary differential triads, as the following triplets:
$\mathcal{T}_{i}:=(\mathbf{A}_{i},d,\Omega_{i})$ (10)
which are the ‘discrete’ analogues of the ADG-theoretic differential triads $(\mathbf{A},\partial,\mathbf{\mathcal{E}}\simeq\Omega)$. [^14] Refer to [38, 39, 44, 53, 54, 55] for more detailed definitions, further interpretational discussion and relevant results.
1.1.6 Sixth encounter with ‘Kinematical’ Functoriality: The Category of Differential Triads
In [64, 65], Papatriantafillou observed that the ADG-theoretic differential triads form a very homologically rich category: the category of differential triads, whose objects are differential triads and whose arrows are differential structure preserving sheaf morphisms.
What is very interesting for us here, as observed in [53, 54, 55, 70, 73], is that there is a contravariant functor between the category of finitary posets and the category of finitary differential triads of incidence algebras. As a result, as the inverse system of Sorkin’s finitary posets was seen to possess an inverse (or projective) limit space, the corresponding, categorically dual, inductive system of finitary incidence algebras was seen to form a direct (or inductive) limit space, consistent and in agreement with Papatriantafillou’s results in [65]. [^15] Subsequently, this observation was crucial in our idea of promoting our category of finitary differential triads into a topos-like structure [37, 70, 73], having (at least finite) categorical limits (direct/inductive) and colimits (inverse/projective).
1.2 Enter ADG: ‘Dynamical’ Functoriality
(Note: Henceforth in this paper, all our finitary considerations, constructions and results presented thus far carry on mutatis mutandis to the general ADG-theoretic constructions. Thus, the finitary case is obtained from the general ADG-theory [38, 39, 44] simply by adjoining a finitarity index-subscript ‘$i$’ to all the ADG-symbols and constructions. [^16] For instance, a general open cover $\mathcal{U}$ of the background topological space $X$ in ADG (:there coined open coordinate gauge of $X$), becomes $\mathcal{U}_{i}$ in our finitary domain of the theory. Similarly, as we shall see next, the finitary version of the Cartan-Kähler differential operator $d$ or the general ADG $\mathbf{A}$-connection operator $\mathcal{D}$, become $d_{i}$ and $\mathcal{D}_{i}$ respectively in our finitary realm without any loss of generality whatsoever [53, 54, 55, 69, 70, 71, 72, 73].)
With the identification of the aforementioned finitary sheaves of quantum causal sets as reticular versions of vector sheaves in ADG, we swiftly moved on to mathematically model dynamical variations thereof.
1.2.1 Briefly revisiting $\mathbf{A}$-connections in ADG
To that end, we readily appreciated that the Cartan-Kähler differential operator in (9) is a special example of an ADG-theoretic connection on the sheaves of ‘differential forms’ that it functorially acts as a sheaf morphism, albeit a flat connection [38, 39, 44, 53, 54, 55].
In order to dynamically vary the quantum causal sets that dwell as (germs of) local sections in the stalks of the aforementioned sheaves $\mathbf{\mathcal{E}}_{i}$, we need to ‘gauge’ [^17] That is, we need to localise and relativise differential changes relative to arbitrary sets of (covering) open gauges $(\mathcal{U}_{i})$ [38, 44, 53, 54, 55]. the flat $d$ to a more general (:curved) connection $\mathcal{D}$,
$d\longrightarrow\mathcal{D}$ (11)
noindent which is also defined functorially as an $\mathbf{A}$-linear, Leibnizian sheaf morphism, acting on the relevant module sheaves as follows:
$\mathcal{D}:~{}\mathbf{\mathcal{E}}\longrightarrow\mathbf{\mathcal{E}}\otimes_{\mathbf{A}}\mathbf{\Omega}\cong\mathbf{\Omega}\otimes_{\mathbf{A}}\mathbf{\mathcal{E}}\equiv\mathbf{\Omega}(\mathbf{\mathcal{E}})$ (12)
With the introduction of $\mathcal{D}$ upon localising or ‘gauging’ the flat differential $d$ relative to a set of local open gauges $\mathcal{U}_{i}$, the latter acquires locally an additional term—the so-called gauge vector field potentials’ term $\mathcal{A}$, [^18] Traditionally, in the Classical Differential Geometry (CDG) $\mathcal{C}^{\infty}$-smooth manifolds $M$ used by Physics [23], the term connection is normally reserved for the so-called gauge vector field potentials $\mathcal{A}_{\mu}^{i}$ (with $\mu$ an external spacetime index, and $i$ an internal gauge symmetry index). On the other hand, from an ADG-theoretic perspective, the denomination connection is a ‘holistic’, ‘unitary’ one, pertaining to $\mathcal{D}$ as a whole, and not referring to its ‘contingent’ local split by a choice of gauge $\mathcal{U}_{i}$ as in (13) above. Read on. as follows:
$d\longrightarrow\mathcal{D}|_{U\in\mathcal{U}_{i}}=d+\mathcal{A}|_{U}$ (13)
As alluded to in the last footnote, from an ADG-theoretic point of view, the connection $\mathcal{D}$ is viewed as a ‘unitary’, autonomous dynamical entity [55, 72, 73], regardless of its local gauge split as in (13). Which brings us to arguably the most important ADG-theoretic definition.
1.2.2 First encounter with ‘Dynamical’ Functoriality: ADG-theoretic Fields
In ADG, a dynamical field is defined as the following pair:
$\mathcal{F}:=(\mathbf{\mathcal{E}},\mathcal{D})$ (14)
That is to say:
A dynamical field is a pair consisting of a vector sheaf $\mathbf{\mathcal{E}}$, localised on an in principle arbitrary $\mathbb{C}$-algebraized space $(X,\mathbf{A})$, and a connection $\mathcal{D}$ acting functorially on its (local) sections as a sheaf morphism.
Four things to highlight here in connection with the fundamental definition above:
  • An ADG-field consists of both the source and the agent of dynamical variability—the connection $\mathcal{D}$, and the recipient of the agent’s dynamical action—the vector sheaf $\mathbf{\mathcal{E}}$, as an autonomous and indivisible/inseparable unit.
  • From a geometric (pre)quantization perspective, the local sections of $\mathbf{\mathcal{E}}$ correspond to quantum particle states [40, 45, 44, 51, 54, 55, 72, 73]. If $\mathbf{\mathcal{E}}$ is a line sheaf $\mathcal{L}$ (:a vector sheaf of rank $n=1$), its local sections represent the quantum particle states of a boson like the ‘photon’, hence the ADG-field $\mathcal{F}_{Max}=(\mathcal{D},\mathcal{L})$ is coined the Maxwell field. More general Yang-Mills ADG-fields are represented as connections $\mathcal{D}$ on vector sheaves $\mathbf{\mathcal{E}}$ of rank $n>1$. ADG-theoretically, we represent them by the pair: $\mathcal{F}_{YM}=(\mathcal{D},\mathbf{\mathcal{E}})$. Finally, the gravitational connections constitute ADG-theoretic Einstein fields: $\mathcal{F}_{Einst}=(\mathcal{D},\mathbf{\mathcal{E}})$.
  • It is important to stress here that, in a very technical and rigorous sense, the vector sheaves $\mathbf{\mathcal{E}}$ correspond to the associated or representation sheaves of the principal group sheaf ${\mathcal{A}}ut(\mathbf{\mathcal{E}})$ of the reversible endomorphisms (:the automorphisms) of $\mathbf{\mathcal{E}}$ [38, 84, 85, 86, 87, 44]. [^19] We recall from [38, 55, 44] that for any vector sheaf $\mathbf{\mathcal{E}}$, ${\mathbf{\mathcal{E}}}nd\mathbf{\mathcal{E}}\equiv{\mathcal{H}}om_{\mathbf{A}}(\mathbf{\mathcal{E}},\mathbf{\mathcal{E}})\cong\mathbf{\mathcal{E}}\otimes_{\mathbf{A}}\mathbf{\mathcal{E}}^{*}=\mathbf{\mathcal{E}}^{*}\otimes_{\mathbf{A}}\mathbf{\mathcal{E}}$, so that ${\mathcal{A}}ut(\mathbf{\mathcal{E}})\simeq\mathbf{\mathcal{E}}nd(\mathbf{\mathcal{E}})^{*}$. It follows that, for a choice of local open gauges $U\in\mathcal{U}_{i}$, ${\mathcal{A}}ut(\mathbf{\mathcal{E}})_{U}\equiv{\mathcal{A}}ut(\mathbf{\mathcal{E}})(U)\simeq\mathbf{\mathcal{E}}nd(\mathbf{\mathcal{E}})^{*}(U)=M_{n}(\mathbf{A}_{U})\equiv M_{n}(\mathbf{A}(U))$, the non-commutative gauge structure group sheaf of $(n\times n)$-matrices having for entries local sections in the structure algebra sheaf $\mathbf{A}$: $\Gamma(U,\mathbf{A})=\mathbf{A}_{U}\equiv\mathbf{A}(U)$. In turn, ${\mathcal{A}}ut(\mathbf{\mathcal{E}})$ is the local relativity and gauge invariance structure group sheaf of the functorial dynamics effectuated by the connection sheaf morphism $\mathcal{D}$ acting dynamically on (the local sections of) $\mathbf{\mathcal{E}}$.
  • The Maxwell $\mathcal{F}_{Max}=(\mathcal{D},\mathcal{L})$, Yang-Mills $\mathcal{F}_{YM}=(\mathcal{D},\mathbf{\mathcal{E}})$ and Einstein $\mathcal{F}_{Einst}=(\mathcal{D},\mathbf{\mathcal{E}})$ ADG-fields can be organised, as objects, into respective categories with categorical sheaf morphisms as arrows between them, coined: the Maxwell Category $\mathcal{T}_{Max}$, the Yang-Mills Category $\mathcal{T}_{YM}$ and the Einstein Category $\mathcal{T}_{Einst}$ categories [44, 49, 51]. In contradistinction to the ‘flat’, ungauged and ‘static-kinematical’ categories of differential triads that we alluded to earlier [64, 65, 44], these three categories are ‘dynamical’ in character, in the sense that object-fields in them obey and satisfy certain dynamical laws of motion [^20] The Maxwell, Yang-Mills and Einstein dynamical (differential) equations, which in turn derive from Lagrangian variation of corresponding action functionals. Read on. and the $\mathbf{A}$-functoriality thereof corresponds to the $\mathbf{A}$-invariance and the ${\mathcal{A}}ut\mathbf{\mathcal{E}}$ local gauge/generalised coordinates’ invariance of the respective dynamical laws of motion [49, 51]. [^21] We shall return to discuss further the categorical implications and the deeper physical interpretation of the generalised ADG-theoretic conception of local gauge invariance as $\mathbf{A}$-invariance and the $\otimes_{\mathbf{A}}$-functoriality of the ADG-theoretic gauge dynamics in the sequel.
    We close this subsection with an important observation regarding the three $ADG$-field categories defined above:
    All three aforementioned ADG-field categories, the Maxwell $\mathcal{T}_{Max}$, the Yang-Mills $\mathcal{T}_{YM}$, and the Einstein category $\mathcal{T}_{Einst}$, are by definition functor categories [37]. [^22] That is to say, the objects in those categories are sheaf morphisms, while the arrows between them are themselves functors. It follows, that if there are functorial correspondences between them and other functor categories, these correspondences will be some kind of natural transformations, and especially, some kind of geometric morphisms [37]. As we will see in the sequel, of special interest to us will be a geometric morphism asssociated with the $\otimes_{\mathbf{A}}$-$\mathrm{Hom}$ adjunction, which effectuates a kind of natural transformation between ADG-field and ADG-curvature space categories. Read on.
1.2.3 Second encounter with ‘Dynamical’ Functoriality: $\mathbf{A}$-invariance is a generalised, functorial form of gauge invariance of the ADG-field dynamics
The discussion above brings us to the all-important issue of gauge covariance, local gauge invariance and their intimate relation to the basic ADG-theoretic notion of $\mathcal{A}$-invariance. [^23] As it has also been observed in past publications [53, 54, 55], in our work we use the symbol $\mathbf{A}$ for the structure sheaf of algebras of generalised ADG-theoretic coordinate functions, as opposed to $\mathcal{A}$ used throughout Mallios’s work [38, 39, 44, 49, 51], as we have reserved the symbol $\mathcal{A}$ for local Einstein gravitational and Yang-Mills local gauge potentials as in (13). In this subsection, we will focus only on [non-abelian] Einstein and Yang-Mills ADG-fields on higher rank vector sheaves, leaving the abelian case of Maxwell fields on line sheaves to their exhaustive treatment in the monograph references [38, 44]. [^24] The epithets abelian and non-abelian above pertain, as in the usual theory [23], to the structure gauge groups being commutative and non-commutative, respectively. Indeed, the principal group sheaves associated with the line sheaves $\mathcal{L}$ of the ADG-theoretic Maxwell fields $\mathcal{F}_{Max}$ above carry abelian unitary (:$\equiv U_{1}(\mathbf{A})$) groups in their stalks, while as we saw couple of footnotes earlier, the structure group sheaf ${\mathcal{A}}ut\mathbf{\mathcal{E}}$ of dynamical automorphisms of the vector sheaf $\mathbf{\mathcal{E}}$ of, say, the Einstein field $\mathcal{F}_{Einst}$ is locally homomorphic to $M_{n}(\mathbf{A}(U))$, which is manifestly noncommutative.
The $\mathbf{A}$-Functoriality of the Curvature of an $\mathbf{A}$-Connection: $R$ is an $\otimes_{\mathbf{A}}$-tensor. To that end, we first recall from [38, 39, 44, 53, 54, 55] the general functorial ADG-theoretic definition of the curvature $R$ of an $\mathbf{A}$-connection $\mathcal{D}$ as the following $\mathbf{A}$)-morphism of $\mathbf{A}$-modules: [^25] With a vector sheaf $\mathbf{\mathcal{E}}$, as explained before, regarded as asheaf of differential $\mathbf{A}$-modules, with structure sheaf $\mathbf{A}$, that is locally isomorphic to $\mathbf{A}^{n}(U)$.
We first define the 1st prolongation of $\mathcal{D}$ to be the following $\mathbf{C}$-linear vector sheaf morphism:
$\mathcal{D}^{1}:~{}\mathbf{\Omega}^{1}(\mathbf{\mathcal{E}})\longrightarrow\mathbf{\Omega}^{2}(\mathbf{\mathcal{E}})$ (15)
satisfying section-wise relative to $\mathcal{D}$:
$\mathcal{D}^{1}(s\otimes t):=s\otimes\text{{d}}t-t\,{\scriptstyle\wedge}\,\mathcal{D}s,~{}(s\in\mathbf{\mathcal{E}}(U),t\in\mathbf{\Omega}^{1}(U),U~{}\mathrm{open~{}in}~{}X)$ (16)
We are now in a position to define the curvature $R$ of an $\mathbf{A}$-connection $\mathcal{D}$ by the following triangular commutative diagram:
${\mathbf{\mathcal{E}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R(\mathcal{D})=\mathcal{D}^{1}\circ\mathcal{D}}$$\scriptstyle{\mathcal{D}}$${\mathbf{\Omega}^{1}(\mathbf{\mathcal{E}})\equiv\mathbf{\mathcal{E}}\otimes_{\mathbf{A}}\mathbf{\Omega}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{D}^{1}}$${\mathbf{\Omega}^{2}(\mathbf{\mathcal{E}})\equiv\mathbf{\mathcal{E}}\otimes_{\mathbf{A}}\mathbf{\Omega}^{2}}$
from which we read directly that:
$R\equiv R(\mathcal{D}):=\mathcal{D}^{1}\circ\mathcal{D}$ (18)
Therefore, any time we have the $\mathbf{C}$-linear morphism $\mathcal{D}$ and its prolongation $\mathcal{D}^{1}$ at our disposal, we can define the curvature $R(\mathcal{D})$ of the connection $\mathcal{D}$ [^26] In connection with (18), one can justify our earlier remark that the standard Cartan-Kähler (exterior) differential operator $d\equiv d^{0}$ is a flat type of connection, since: $R(d)=d\circ d\equiv d^{2}=0$, which is secured by the well known nilpotency of the usual Cartan-Kähler (exterior) differential operator $d$ [23, 38, 44, 55]). In a (co)homological-algebraic sense, the curvature of an algebraic connection measures the ‘obstruction’ to or the ‘deviation’ from the nilpotency of the connection (:differential) [53, 54, 55].
As a matter of fact, it is rather straightforward to see that, for $\mathbf{\mathcal{E}}$ a vector sheaf, $R(\mathcal{D})$ is functorially defined as an $\mathbf{A}$-morphism of $\mathbf{A}$-modules, in the following sense:
$\begin{array}[]{c}R\in{\mathrm{Hom}}_{\mathbf{A}}(\mathbf{\mathcal{E}},\mathbf{\Omega}^{2}(\mathbf{\mathcal{E}}))={\mathcal{H}}om_{\mathbf{A}}(\mathbf{\mathcal{E}},\mathbf{\Omega}^{2}(\mathbf{\mathcal{E}}))(X)\cr\mathbf{\Omega}^{2}({\mathbf{\mathcal{E}}}nd\mathbf{\mathcal{E}})(X)=Z^{0}(\mathcal{U},\mathbf{\Omega}^{2}({\mathbf{\mathcal{E}}}nd\mathbf{\mathcal{E}}))\end{array}$ (19)
where, as usual, $\mathcal{U}_{i}=\{U_{\alpha}\}_{\alpha\in I}$ is an open cover of the base topological space $X$ and $Z^{0}(\mathcal{U},\mathbf{\Omega}^{2}(\mathcal{E}nd\mathbf{\mathcal{E}}))$ the $\mathbf{A}(U)$-module of 00-cocycles of $\mathbf{\Omega}^{2}(\mathbf{\mathcal{E}}nd\mathbf{\mathcal{E}})$ relative to the $\mathcal{U}_{i}$-covering of $X$ [^27] One may wish to recall again that, for a vector sheaf $\mathbf{\mathcal{E}}$ like the one involved in (19) above: ${\mathbf{\mathcal{E}}}nd\mathbf{\mathcal{E}}\equiv{\mathcal{H}}om_{\mathbf{A}}(\mathbf{\mathcal{E}},\mathbf{\mathcal{E}})\cong\mathbf{\mathcal{E}}\otimes_{\mathbf{A}}\mathbf{\mathcal{E}}^{*}=\mathbf{\mathcal{E}}^{*}\otimes_{\mathbf{A}}\mathbf{\mathcal{E}}$.
1.2.4 The Non-Tensorial and the Tensorial Character of $\mathcal{D}$ and $R$, respectively, under Local ‘Gauge-Coordinate’ Transformations
(Note: In the following presentation and discussion, we are not going to specify what ADG-connections and their curvatures we are talking about. The reader can assume that the connections are either Einstein-Lorentzian or Yang-Mills, in the sense that the arguments below apply mutatis mutandis to both.)
In this subsection, we recall from [38, 55, 44] a very subtle and important for our arguments in the sequel ADG-theoretic result, which may be distilled down to the following two statements:
$\bullet$ The ADG-theoretic connection $\mathcal{D}$ is only a $\mathbf{C}$-linear sheaf morphism (hence not an $\otimes_{\mathbf{A}}$-tensor); [^28] Where $\mathbf{C}$ is just the constant sheaf of $\mathbb{C}$-numbers [38, 55, 44]. while;
$\bullet$ The ADG-theoretic curvature $R$ is a full $\mathbf{A}$-structure sheaf morphism (hence a pure $\otimes_{\mathbf{A}}$-tensor). [^29] With $\otimes_{\mathbf{A}}$ the homological algebraic (:categorical) tensor product functor. In the sequel, we will return to explain and discuss in more detail the paramount importance of $\otimes_{\mathbf{A}}$ and its adjoint functor $\mathrm{Hom}$ [35, 36, 37] for the dynamical functoriality of our (finitary) ADG-perspective on gravity and gauge theories, and its cogent physical interpretation in the quantum deep.
Two equivalent statements to the ones above, which the theoretical physicist/mathematician who is familiar with the usual differential geometry of gauge theory, which employs smooth fiber bundles over a $\mathcal{C}^{\infty}$-smooth differential spacetime manifold $M$ can straightforwardly understand [23], are the following:
$\bullet$ $\mathcal{D}$ does transforms inhomogeneously (:non-tensorially or affinely) under local gauge-coordinate transformations; while;
$\bullet$ $R$ does indeed transform homogeneously (:tensorially) under local gauge-coordinate transformations.
To explicate in detail what the above statements mean, let us recall briefly from [38, 55, 44] how the ADG-connection $\mathcal{D}$ and the ADG-curvature $R$ behave (:transform) respectively under local general coordinate-gauge changes.
1.2.5 Local Gauge Transformation of $\mathcal{D}$
[…]
$f_{j}=\sum_{i=1}^{n}g_{ij}e_{i}$ (20)
which, plainly, is a local (i.e., relative to $U\cap V$) section of the ‘natural’ structure group sheaf $\mathcal{G}\mathcal{L}(n,\mathbf{A})$ of $\mathbf{\mathcal{E}}$ [^32] As noted earlier, one may recognise $\mathcal{G}\mathcal{L}(n,\mathbf{A})$ above as the local version of the automorphism principal group sheaf ${\mathcal{A}}ut\mathbf{\mathcal{E}}$ of $\mathbf{\mathcal{E}}$. The adjective ‘local’ here pertains to the fact mentioned earlier that ADG assumes that $\mathbf{\mathcal{E}}$ is locally isomorphic to $\mathbf{A}^{n}$.—that is to say, $g_{ij}\in\mathrm{GL}(n,\mathbf{A}(U\cap V))=\mathcal{G}\mathcal{L}(n,\mathbf{A})(U\cap V)$.
Without going into the details of the derivation, which can be found in [38, 39, 44], we note that under such a local gauge transformation $g$, the gauge potential part $\mathcal{A}$ of $\mathcal{D}$ in (13) transforms as follows:
$\mathcal{A}^{{}^{\prime}}=g^{-1}\mathcal{A}g+g^{-1}\partial g$ (21)
a way we are familiar with from the usual differential geometry of the smooth fiber bundles of gauge theories [23]. For completeness, it must be noted here that, in (21), $\mathcal{A}\equiv(\mathcal{A}_{ij})\in M_{n}(\mathbf{\Omega}^{1}(U))=M_{n}(\mathbf{\Omega}^{1})(U)$ and $\mathcal{A}^{{}^{\prime}}\equiv(\mathcal{A}^{{}^{\prime}}_{ij})\in M_{n}(\mathbf{\Omega}^{1}(V))=M_{n}(\mathbf{\Omega}^{1})(V)$.
The transformation of $\mathcal{A}$ under local gauge changes is called inhomogeneous, non-tensorial or affine in the usual gauge-theoretic parlance [23] precisely because of the (additional to the homogeneous) term $g^{-1}\partial g$.
1.2.6 Local Gauge Transformation of $R$
On the other hand, we read directly from [38, 55, 44] that under similar local gauge-coordinate changes, the curvature $R(\mathcal{D})$ of the ADG-connection $\mathcal{D}$ transforms purely homogeneously or tensorially, as follows:
To that end, let again $g\equiv g_{ij}\in\mathcal{G}\mathcal{L}(n,\mathbf{A})(U\cap V)$ be the change-of-gauge matrix we considered in (20) in connection with the transformation law of gauge potentials $\mathcal{A}_{ij}$. Again, without going into the technical details of the derivation, we bring forth from [38, 55, 44] the following local transformation law of gauge field strengths:
$\begin{array}[]{c}\mathrm{for~{}a~{}local~{}frame~{}change:~{}}e^{U}\stackrel{{\scriptstyle g}}{{\longrightarrow}}e^{V}(U,V\in\mathcal{U}{i}~{}\mathrm{covering}~{}X),\cr\mathrm{the~{}curvature~{}transforms~{}as:}~{}R\stackrel{{\scriptstyle g}}{{\longrightarrow}}R^{{}^{\prime}}=g^{-1}Rg\end{array}$ (22)
the form of which we are familiar with from the usual differential geometric (i.e., smooth fiber bundle-theoretic) treatment of gauge theories [23]. For completeness, we remind ourselves here that, in (21) above, $R^{U\cap V}\equiv(R^{U\cap V}_{ij})\in M_{n}(\mathbf{\Omega}^{2}(U\cap V))$—an $(n\times n)$-matrix of sections of local $2$-forms in $\mathbf{\Omega}^{2}$.
The transformation of $R$ under local gauge-coordinate changes is called homogeneous, tensorial or covariant in the usual smooth fiber bundle gauge-theoretic parlance [23].
As a last important observation before we move on to explicate the $\mathbf{A}$-invariant and its associated $\otimes_{\mathbf{A}}$-functorial character of the dynamical equations of motion from Einstein gravity and free Yang-Mills gauge theories, we note:
1.2.7 ADG-curvature spaces and ADG-curvature field categories
As we increase by a notch the level of abstraction and generality, from [38, 55, 44] we note the definition of ADG-curvature spaces as the following quintuples:
$(\mathbf{A},\mathcal{D},\mathbf{\Omega}^{1},\mathcal{D}^{2},\mathbf{\Omega}^{2})$ (23)
consisting of $\mathbf{A}$-modules and $\mathbf{C}$-linear morphisms between them, which, in turn, by the very definition of the ADG-curvature field in (17), reduce to the following duet representing the ADG-curvature fields:
$(\mathbf{A},R(\mathcal{D}))$ (24)
1.2.8 Third encounter with ‘Dynamical’ Functoriality: three ADG-curvature field functor categories
In much the same way that we defined earlier the three functor categories of ADG-connection fields: the Maxwell $\mathcal{T}_{Max}$, the Yang-Mills $\mathcal{T}_{YM}$, and the Einstein category $\mathcal{T}_{Einst}$ earlier, we can similarly define here:
[…]
1.2.9 Lagrangean Action Derivation of Vacuum Einstein Gravity and Free Yang-Mills Theories
Now that we have recalled the essential characteristics and local gauge transformation behaviour of the affine ADG-connections and their curvature $\otimes_{\mathbf{A}}$-tensors, we note that the ADG-theoretic versions of the dynamical free Yang-Mills and vacuum Einstein equations both derive from (the variation of) respective Yang-Mills ($\mathcal{Y}\mathcal{M}$) and Einstein-Hilbert ($\mathcal{E}\mathcal{H}$) Lagrangean action functionals, as follows: [^34] For the equations above, see [38, 44] for technical definitions and details.
$\begin{array}[]{c}\mathcal{E}\mathcal{H}_{\mathbf{\mathcal{E}}}(\mathcal{D})=\int_{X}tr(R_{Ric}(\mathcal{D}))\;\stackrel{{\scriptstyle\delta\mathcal{A}}}{{\longrightarrow}}\;R_{Einst}(\mathbf{\mathcal{E}})=0\cr\cr\mathcal{Y}\mathcal{M}_{\mathbf{\mathcal{E}}}(\mathcal{D})=\frac{1}{2}\int_{X}tr(R_{YM}\wedge\star R_{YM})\;\stackrel{{\scriptstyle\delta\mathcal{A}}}{{\longrightarrow}}\;\Delta_{\mathbf{\mathcal{E}}nd\mathbf{\mathcal{E}}}^{2}(R_{YM})=0\end{array}$ (25)
in which we read from [38, 55, 44] that $R_{Einst}$ is the Ricci Scalar [^35] For expository completeness, we briefly recall from [55] that given a (real) Lorentzian vector sheaf $(\mathbf{\mathcal{E}},\rho)$ of rank $n$ equipped with a non-flat Lorentzian $\rho$-metric $\mathbf{A}$-connection $\mathcal{D}$, one can define the following Ricci curvature operator $R_{Ric}$ relative to a local gauge $U\in\mathcal{U}_{i}$ of $\mathbf{\mathcal{E}}$ : $R_{Ric}(\,.\,,\,\!s)t\in(\mathbf{\mathcal{E}}nd\mathbf{\mathcal{E}})(U)=M_{n}(\mathbf{A}(U))$, for local sections $s$ and $t$ of $\mathbf{\mathcal{E}}$ in $\mathbf{\mathcal{E}}(U)=\mathbf{A}^{n}(U)=\mathbf{A}(U)^{n}$. Thus, the Ricci curvature here $R_{Ric}$ is an $\mathbf{\mathcal{E}}nd\mathbf{\mathcal{E}}$-valued operator, a curvature endomorphism of $\mathbf{\mathcal{E}}$. Moreover, since $R_{Ric}$ is matrix-valued, one can take its trace, thus define the following Ricci scalar curvature operator $R_{Einst}:=tr(R_{Ric}(\,.\,,\,\!s)t)$, which, plainly, is $\mathbf{A}(U)$-valued. of the ADG-theoretic Einstein-Lorentzian metric connection field $\mathcal{D}$, while $R_{YM}$ is the Yang-Mills gauge field strength of the homonymous ADG-theoretic Yang-Mills connection field $\mathcal{D}$.
1.3 Miscellaneous Remarks on ‘Dynamical Functoriality’ and $\mathbf{A}$-Invariance-cum-Covariance of the ADG-Field Autodynamics: A Unified, Pure Gauge, Smooth Base Spacetime Manifoldless and Finitistic ADG-Theoretic Quantum Field ‘Solipsism’ of the 3rd Kind
In this subsection we make eight conceptual and technical remarks on the ADG-theoretic perspective on the Einstein and Yang-Mills field dynamics in (25) above. We itemise our remarks as follows:
1.3.1 ADG-Kinematics: The Affine Space of $\mathbf{A}$-Connections
From the vacuum Einstein and the free Yang-Mills dynamical equations à-la ADG in (25) above, it follows that the sole dynamical variable in our ADG-theoretic perspective on gravity and gauge theories is the local $\mathbf{A}$-connection $\mathcal{D}$. That is, as equation (25) depicts above, the dynamical equations derive from the variation ($\delta\mathcal{A}$) of the Einstein-Hilbert and Yang-Mills Lagrangian action functionals with respect to the local gauge potential part ($\mathcal{A}$) of the gravitational and Yang-Mills $\mathbf{A}$-connections $\mathcal{D}$ on their respective vector sheaves $\mathbf{\mathcal{E}}$.
Thus, as emphasised in [55], the sole dynamical variable in our ADG-theoresis of vacuum Einstein gravity and free Yang-Mills theories is the ADG-theoretic Einstein connection field pairs $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$ and $\mathcal{F}_{YM}=(\mathbf{\mathcal{E}},\mathcal{D})$ defined by (14) earlier, within their respective categories $\mathcal{T}_{Einst}$ and $\mathcal{T}_{YM}$.
As also highlighted in [55], it follows that the generalised ‘kinematical’ space of the theory is the affine space $\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})$ of $\mathbf{A}$-connections $\mathcal{A}$ on $\mathbf{\mathcal{E}}$. Moreover, since the Lagrangians involved in (25) are invariant under the group sheaf $\mathcal{G}(\mathbf{\mathcal{E}})={\mathcal{A}}ut\mathbf{\mathcal{E}}$ of local automorphisms (:local gauge transformations) of $\mathbf{\mathcal{E}}$,
the relevant kinematical space is the moduli space $\textsf{A}/\mathcal{G}=\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut\mathbf{\mathcal{E}}$ of gauge equivalent $\mathbf{A}$-connections $\mathcal{A}$ on $\mathbf{\mathcal{E}}$ [^36] Recall from [55] that $\mathcal{G}(\mathbf{\mathcal{E}})|_{U}={\mathcal{A}}ut\mathbf{\mathcal{E}}|_{U}=(\mathrm{Aut}_{\mathbf{A}}\mathbf{\mathcal{E}})(U):=\Gamma(U,{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}})\equiv M^{n}_{U}(\mathbf{A})$. Thus, $\textsf{A}/\mathcal{G}=\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut\mathbf{\mathcal{E}}$ is the so-called $\mathcal{G}$-orbit space as the structure gauge group sheaf ${\mathcal{A}}ut\mathbf{\mathcal{E}}(\mathbf{A})$ cuts through the affine space $\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})$, carving out ‘paths’ or ‘orbits’ of gauge equivalent connections in the process, which, in turn, leave the corresponding curvature Lagrangians in (25) invariant under (local) gauge transformations.
Hence the integration sign in the dynamical action functionals in (25), which supposedly extends over the base topological space $X$, in effect extends over the moduli space $\textsf{A}/\mathcal{G}=\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut\mathbf{\mathcal{E}}$ of gauge equivalent connections. [^37] From [55, 49, 51, 44] we read that a suitably defined ADG-theoretic Radon-type of $\mathbf{A}$-linear continuous integration measure $d\mu$ on a suitably topologised $\textsf{A}/\mathcal{G}$ ($d\mu:\;\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut\mathbf{\mathcal{E}}\longrightarrow\mathbf{A}$) is expected to render rigorous the dynamical action integrals in (25).
1.3.2 $\mathbf{A}$-Invariance and ‘Dynamical’ Functoriality
We noted earlier a fundamental difference in ADG between an $\mathbf{A}$-connection $\mathcal{D}$ and its curvature $R(\mathcal{D})$, namely that,
The curvature $R(\mathcal{D})$ is an $\otimes_{\mathbf{A}}$-tensor, while the connection $\mathcal{D}$ itself is not.
In other words,
The curvature $R(\mathcal{D})$ respects our (algebras) of generalised measurements in $\mathbf{A}$, while the connection $\mathcal{D}$ itself does not—it ‘eludes’ them.
In still equivalent parlance,
The curvature $R(\mathcal{D})$ is a ‘geometrical object’, while the connection $\mathcal{D}$ itself is an ‘algebraic object’. [^38] We will make this statement mathematically much more precise and rigorous in the sequel when we discuss the fundamental $\otimes_{\mathbf{A}}-\mathrm{Hom}$-adjunction ‘geometric morphism equivalence’ between the ADG Einstein and/or Yang-Mills connection field and curvature field functor categories $\mathcal{T}_{Einst/YM}$ and $\mathcal{C}_{Einst/YM}$, respectively.
Yet,
the dynamics on $\mathbf{\mathcal{E}}$, generated by the connection field $\mathcal{D}$ acting on $\mathbf{\mathcal{E}}$ and expressed as a differential equation on it as in (25), is derived from an action principle involving the curvature of the connection.
As such,
the dynamical equations of motion on $\mathbf{\mathcal{E}}$, which are derived from an action principle involving the curvature of the connection, is gauge invariant, hence our free gauge choices of generalised coordinate measurements in $\mathbf{A}$ respect the dynamics, and the physical laws are independent of our measurements in $\mathbf{A}$.
Thus, in toto,
The physical laws are $\mathbf{A}$-invariant.
Which brings us to a fundamental observation, in connection with Utiyama’s Theorem, that we read directly from [51].
1.3.3 The Algebra-Geometry Duality: The $\mathrm{Hom}_{\mathbf{A}}-\otimes_{\mathbf{A}}$ Functorial Adjunction between the ADG-Field and Curvature Categories
Below, we quote Mallios verbatim from [51]:
“…Utiyama’s theorem, relates/characterizes the ‘$\mathbf{A}$-invariance’ of what we may call $\mathbf{A}$-connection Lagrangian through that one of the corresponding curvature Lagrangian. So the aforementioned two notions (of ‘Lagrangians’) are, in effect (physically) equivalent, through/due to the ‘$\mathbf{A}$-invariance’…” [^39] Throughout this quotation we have been faithful to the emphasis placed by Mallios on certain key words in the original paper [51].
which leads us to Mallios’s telling remarks of what he calls The Fundamental (Physical) Adjunction: [^40] Again quoting Mallios verbatim from [51].
“…The Fundamental (Physical) Adjunction: Thus, the basic Homological (:categorical) $Hom-\otimes$ adjunction, corresponds, within the context of ADG, to the fundamental physical adjunction, effectuated by the following ‘adjoint pair of functors’:
$\mathbf{A}-connection\;(:field,\;`potential^{\prime})\leftrightarrows\;curvature\;(:`field\;strength^{\prime})$
The above can actually be perceived, as describing the whole function of a physical law, hence, in fact, of the Nature herself…” [^41] Again,throughout the quotation above we have been faithful to the emphasis placed by Mallios on certain key words in the original paper [51].
And Mallios concludes Section 2 of the paper [51] with the following intuitively telling paragraph:
“…On the other hand, the connecting function of a given adjunction, is in effect a natural transformation of functors. Consequently, the latter should still preserve ‘$\mathbf{A}$-invariance’ of the adjunction, with respect to any $\mathbf{A}$-invariant function’, referring to any one of the two associated functors through the adjunction: One gets at it, just, based on the very definitions [^42] Given before in the paper [51]. and on the ‘functorial nature’ of ADG…” [^43] Once again, throughout the quotation above, we have been faithful to the emphasis placed by the author on certain key words in the original text [51].
Now, in view of our presentation and arguments in the present paper, we are in a position to distill and further mathematically formalise and explicate Mallios’s remarks above on functoriality, adjunction and natural transformation of functors.
1.3.4 Fourth Encounter with ‘Dynamical’ Functoriality: Mallios’s Fundamental (Physical) Adjunction Explicated and Interpreted
The categorical adjunction between the connection $\mathcal{D}$ and its curvature $R(\mathcal{D})$
$\mathbf{A}-connection\rightleftarrows\mathbf{A}-curvature$ (26)
that Mallios emphasises in the excerpt from [51] in connection with Utiyama’s Theorem displayed above, can now be cast in a mathematically rigorous and precise functorial form, as a functorial correspondence between the respective categories of Einstein (or Yang-Mills) connection fields $\mathcal{T}_{Einst}$ (or $\mathcal{T}_{YM}$) and their corresponding categories of Einstein (or Yang-Mills) curvature field strengths $\mathcal{C}_{Einst}$ (or $\mathcal{C}_{YM}$), as follows:
[…]
with the pair of ‘opposite direction maps’ $(\otimes_{\mathbf{A}},\mathrm{Hom})$ above corresponding to the fundamental homological algebraic adjunction that Mallios alludes to.
More technically speaking, $\otimes_{\mathbf{A}}$ is the homological (left-adjoint) tensor product functor [^44] In the category of $\mathbf{A}$-modules that the vector sheaves $\mathbf{\mathcal{E}}$ of ADG belong. and $\mathrm{Hom}$ is its right-ajoint functor [37]. When paired together, the pair:
$\mathcal{G}\mathcal{M}_{\mathbf{A}}:=(\otimes_{\mathbf{A}},\mathrm{Hom}_{\mathbf{A}})$ (28)
[…]
At the same time, the fact that the pair $(\otimes_{\mathbf{A}},\mathrm{Hom})$ indeed constitutes a categorical adjunction, derives directly from the way $\otimes_{\mathbf{A}}$ and $\mathrm{Hom}_{\mathbf{A}}$ act on the corresponding categories. Again, in the general case, we read from [37] that $-\otimes X$ is the left-adjoint (functor) and $\mathrm{Hom}(X,-)$ the right-ajoint (functor), because they act as a pair of maps as follows:
$\mathrm{Hom}(Y\otimes X,Z)\simeq\mathrm{Hom}(Y,\mathrm{Hom}(X,Z))$ (29)
Thus, in view of the general definition of the action of an adjunction as in (29) above, we are now in a position to see directly that, indeed:
In ADG, the curvature $R(\mathcal{D})$ of a connection is the $\otimes_{\mathbf{A}}$-morph (:image) of its connection.
which we can verify directly from the curvature’s definition in terms of the action of the $\otimes_{\mathbf{A}}$ and $\mathrm{Hom}_{\mathbf{A}}$ functors in (17) and (19) earlier:
$\begin{array}[]{c}R:\;\mathbf{\mathcal{E}}\stackrel{{\scriptstyle\mathrm{Hom}_{\mathbf{A}}}}{{\longrightarrow}}\mathbf{\Omega}^{2}(\mathbf{\mathcal{E}})\equiv\mathbf{\mathcal{E}}\otimes_{\mathbf{A}}\mathbf{\Omega}^{2}\cr R\in{\mathrm{Hom}}_{\mathbf{A}}(\mathbf{\mathcal{E}},\mathbf{\Omega}^{2}(\mathbf{\mathcal{E}}))={\mathcal{H}}om_{\mathbf{A}}(\mathbf{\mathcal{E}},\mathbf{\Omega}^{2}(\mathbf{\mathcal{E}}))(X)\end{array}$ (30)
Recalling again from footnote 26 that, for any ADG-theoretic vector sheaf $\mathbf{\mathcal{E}}$ like the one involved in (30) above: ${\mathbf{\mathcal{E}}}nd\mathbf{\mathcal{E}}\equiv{\mathcal{H}}om_{\mathbf{A}}(\mathbf{\mathcal{E}},\mathbf{\mathcal{E}})\cong\mathbf{\mathcal{E}}\otimes_{\mathbf{A}}\mathbf{\mathcal{E}}^{*}=\mathbf{\mathcal{E}}^{*}\otimes_{\mathbf{A}}\mathbf{\mathcal{E}}$.
We can distill all the above discussion and express the geometric morphism functorial equivalence between the connection-field and curvature-field Einstein and/or Yang-Mills functor categories $\mathcal{T}_{Einst/YM}$ and $\mathcal{C}_{Einst/YM}$ depicted in equations (27) and (28) above, as follows:
[…]
with the first two lines in (31) above, reading:
The curvature (Einstein and/or Yang-Mills field categories) are the geometric morphs (:images) of the corresponding connection (field categories).
while the map in the third line of (31) above can be interpreted as stipulating that:
The (Einstein and/or Yang-Mills) curvature field is the $\otimes_{\mathbf{A}}$-image of its connection field.
which in turn vindicates what we established earlier, namely that:
The curvature is a geometrical, $\mathbf{A}$-tensorial object (:an $\otimes_{\mathbf{A}}$-tensor or an $\mathbf{A}$-invariant morphism), while its connection is not.
This gives a raison d’être and vindicates the epithet geometric in the geometric morphism denomination of the pair of maps $\mathcal{G}\mathcal{M}_{\mathbf{A}}:=(\otimes_{\mathbf{A}},\mathrm{Hom}_{\mathbf{A}})$ above.
1.4 $\mathcal{G}\mathcal{M}$-Dynamical Functoriality, Mallios’s $\mathbf{A}$-Invariance and Gauge Invariance of ADG-GT
Now that we have explicated the subtle and technical sense in which a geometric morphism adjunction $\mathcal{G}\mathcal{M}$ links the ADG-theoretic Einstein/Yang-Mills connection field category $\mathcal{T}_{Einst/YM}$ with its ADG-theoretic Einstein/Yang-Mills curvature field ‘counterpart-equivalent’ category $\mathcal{T}_{Einst/YM}$, we are in a position to further support Mallios’s remarks on $\mathbf{A}$-invariance in connection with gauge invariance in [51].
To this end, we quote verbatim from Section 3 of [51] the displayed paragraph before Theorem 3.1:
“…Therefore, one thus realizes that, the fundamental adjunction [^46] Our geometric morphism $\mathcal{G}\mathcal{M}$ in equations (27) and (28) earlier., preserves the $\mathbf{A}$-invariance, for any $\mathbf{A}$-invariant function, [^47] Especially, for the $\mathbf{A}$-invariant Einstein and Yang-Mills action functionals in (25) above, which, as we saw earlier, are $\mathbf{A}$-valued functionals defined on the moduli space $\textsf{A}_{\mathbf{A}}/\mathcal{G}=\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut\mathbf{\mathcal{E}}$ of gauge equivalent $\mathbf{A}$-connections on $\mathbf{\mathcal{E}}$. pertaining to the two basic functors appearing in the aforesaid adjunction…” [^48] That is, the homological $\mathrm{Hom}_{\mathbf{A}}$ and $\otimes_{\mathbf{A}}$ adjoint functors constituting $\mathcal{G}\mathcal{M}_{\mathbf{A}}$ in equations (27) and (28) earlier.
which in turn leads to the following central result (:Theorem 3.1) in [51], coined Utiyama’s Principle therein: [^49] Again, quoted exactly as it appears in [51].
“…Theorem 3.1 Any ‘gauge invariance’ of an appropriate ‘Lagrangian’ for ($\mathbf{A}$-)connections is equivalent to a similar invariance of the corresponding Lagrangian for the associated curvature with the ($\mathbf{A}$-)connection at issue…
Thus, we are now in a position to distill and re-express the deep relation between our notion of ‘Dynamical’ Functoriality, with Mallios’s $\mathbf{A}$-invariance and the structure group ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$-gauge invariance (of the dynamical action functionals) of our ADG-theoresis on Vacuum Einstein Gravity and Free Yang-Mills Theories in (25), as follows:
Fundamental Theorem of ADG-Gauge Theory. The ‘dynamical’ geometric morphism $\mathcal{G}\mathcal{M}_{\mathbf{A}}$ preserves $\mathbf{A}$-invariance and entails the structure group sheaf $\mathcal{G}\equiv{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$-invariance of the ADG-theoretic dynamical equations of motion for Vacuum Einstein Gravity and Free Yang-Mills Theory.
In view of our physical interpretation of $\mathbf{A}$ earlier in this paper and throughout our work on ADG-GT [53, 54, 55, 44, 69, 70, 71, 72, 73] as the algebra (sheaf) of our generalised coordinate measurements, [^50] That is, our localised and gauged $\mathbf{A}$-valued measurements based on an open cover $\{U\}=\mathcal{U}_{i}$ of local open sets $U$ of the base topological space $X$ employed by the theory. an important Corollary to the Fundamental Theorem above goes the other way around, as follows:
Corollary to the Fundamental Theorem of ADG-Gauge Theory. The algebra (sheaf) of our generalised coordinate measurements $\mathbf{A}$ respects (i.e. it is ‘non-perturbing’ and it leaves invariant) the functorial ADG-theoretic gauge field dynamics for Vacuum Einstein Gravity and Free Yang-Mills theories, hence, in return, it entails and almost ‘mandates’ that the (local) Relativity Group (sheaf) of the theory is $\mathcal{G}\equiv{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ [^51] That the local relativity group in ADG-GT is ‘naturally’ ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ (:see subsection next) has been amply expounded in [53, 54, 55, 44, 49, 50, 69, 70, 71, 72, 73]. Thus, $\mathbf{A}$-invariance, via the ‘dynamical’ geometric morphism $\mathcal{G}\mathcal{M}_{\mathbf{A}}$, which in turn entails the $\mathcal{G}={\mathcal{A}}ut\mathbf{\mathcal{E}}$-gauge invariance of the ADG-field dynamics (expressed via the geometric morph of the connection field—the field’s curvature), corresponds to a dynamical version of the Kleinian conception of geometry. [^52] According to Felix Klein, ‘the geometry of an object is all the transformations of it that leave it invariant[32].
The discussion above brings us to an important, yet subtle, technical and interpretational matter of ADG-GT.
1.4.1 Note on the Natural Transformation character of the Geometric Morphism $\mathcal{GM}_{\mathbf{A}}$: the ‘Naturality’ of the Functorial Dynamics of ADG-GT
Since, as we alluded to numerous times throughout this paper, the ADG-theoretic connection and curvature field categories are functor categories [51, 49, 50, 44], the two adjoint functors constituting the geometric morphism $\mathcal{G}\mathcal{M}_{\mathbf{A}}:=(\otimes_{\mathbf{A}},\mathrm{Hom}_{\mathbf{A}})$ above are examples of natural transformations [37]. [^53] In a nutshell, and quite heuristically, a natural transformation $\mathcal{N}$ connects or maps one particular functor $\mathcal{F}:\,A\rightarrow B$ to another particular functor $\mathcal{G}:\,A\rightarrow B$ between two categories $A$ and $B$. At the same time, $\mathcal{N}$ does not need to apply to every functor in some category of functors [37].
This, too, was prophetically anticipated by Mallios in [51]: [^54] Excerpt from quotation earlier.
“…On the other hand, the connecting function of a given adjunction, is in effect a natural transformation of functors. Consequently, the latter should still preserve ‘$\mathbf{A}$-invariance’ of the adjunction, with respect to any $\mathbf{A}$-invariant function’, referring to any one of the two associated functors through the adjunction…”
The fitting physico-mathematical ‘pun’ here is that:
The $\mathbf{A}$-invariant and, in extenso, ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$-invariant functorial gauge field dynamical changes in ADG-GT are, categorically-cum-physically speaking, Natural Transformations of the ‘dynamically equivalent’ gauge connection and curvature fields involved therein, via the Natural Transformation Geometric Morphism $\mathcal{G}\mathcal{M}_{\mathbf{A}}:=(\otimes_{\mathbf{A}},\mathrm{Hom}_{\mathbf{A}})$ that interlinks them.
2 Intermezzo: Lateral Technical and Philosophical Offshoots and Repercussions of Functoriality
In this intermediate section, we give very brief accounts and we express them in the form of ‘Aphorisms’, borrowed from previous works, of various technical, conceptual and interpretational-cum-philosophical corollaries and didactics that follow, in one way or another, from both the ’kinematical’ and ‘dynamical’ functoriality of our ADG-theoretic perspective on Finitary Vacuum Einstein-Lorentzian Gravity and Free Yang-Mills Theories, as expounded above.
2.1 Third Gauge Auto-Gravitodynamics from Background Spacetime Manifoldlessness: Gauge Field Solipsism
The quintessential feature of ADG, especially vis-à-vis its novel conceptual import and potential technical applications to Quantum Gravity research [38, 39, 44, 40, 45, 42, 46, 48, 47, 52, 51, 49, 50, 53, 54, 55, 56, 57, 58, 59, 60, 61, 67, 68, 69, 70, 71, 72, 73, 74, 75, 92, 93], is arguably the following:
Aphorism 1: Background Spacetime Manifoldlessness. Mallios’s Abstract Differential Geometry is a purely homological-algebraic (:sheaf and category-theoretic) way of doing and applying ‘Differential Calculus’, with all its technical and conceptual panoply, to many current research fronts in Theoretical and Mathematical Physics such as Quantum Gauge Theories of Matter and Quantum Gravity, but in the manifest absence of a background geometrical $\mathcal{C}^{\infty}$-smooth base (spacetime) manifold.
The deep and wide spectrum of potential import of such background differential spacetime manifoldlessness, especially in Quantum Gauge Theory and Quantum Gravity research, has been expounded in detail over the last two decades in numerous works [40, 41, 45, 46, 49, 50, 44, 53, 54, 55, 69, 70, 71, 72, 73, 60, 61, 92, 93].
One important feature of such a background spacetime manifold independence is that:
Aphorism 2: Dynamical Connection Gauge Field Solipsism. ADG enables us to formulate Vacuum Einstein Gravity and Free Yang-Mills theories as pure gauge theories of the third kind, [^55] We recall directly from [55, 72, 73] that Gauge Theory of the First Kind is Hermann Weyl’s original Global ($U(1)$) Scale Theory of the Electromagnetic Field [88]; Gauge Theory of the Second Kind pertains to the usual current abelian and non-abelian Yang-Mills gauge theories of matter that are localised and gauged based on an externally prescribed and fixed $\mathcal{C}^{\infty}$-smooth base spacetime continuum; while Gauge Theory of the Third Kind is our ADG-theoresis, which is ‘field monic, solipsistic and autonomous’. Read on. in the sense that the sole dynamical variable in the theory is an entirely homologically-algebraically defined $\mathbf{A}$-connection acting on (the sections of) a vector sheaf $\mathbf{\mathcal{E}}$, without any recourse to or dependence on an external background geometrical differential (:$\mathcal{C}^{\infty}$-smooth) spacetime manifold [55, 72, 73, 44, 61]. The ADG-theoretic Gauge Theory of the Third Kind, which regards the $\mathbf{A}$ connection field $(\mathbf{\mathcal{E}},\mathcal{D})$ as the sole dynamical variable, has been coined Half-Order Formalism[55, 72, 73]. [^56] To distinguish it from the original Second Order Formalism of Einstein, involving the smooth spacetime metric $g_{\mu\nu}$ on a background differential spacetime manifold [63], as well as from the more recent First Order Formalism of Ashtekar et al., which, apart from a Lorentzian spin-connection $\mathcal{A}_{\mu}$, it includes the metric-vierbein $e^{a}_{\mu}$ as joint gravitational dynamical variables in the theory [2, 3]. In ADG-GT, dynamics concerns and derives solely from the stalks (of the sheaves involved), not from the base topological space $X$ itself, which is only used for the sheaf-theoretic localisation (and continuous variation) of the ‘geometrical objects’ (:the algebraic connection fields) that live on that surrogate external base space. [^57] To use a Wittgensteinian metaphor here: “Once one climbs up the ladder, one throws the ladder away[91]. Similarly in our ADG-theoresis of gauge theory and gravity, once we have employed the base topological space $X$ as a surrogate scaffolding for the sheaf-theoretic localisation, gauging and continuous dynamical variation of the (Einstein and Yang-Mills) gauge connection fields on it, we formulate the gauge invariant dynamics homologically-algebraically and functorially as equations involving sheaf morphisms as we showed earlier, and we then completely forget about $X$, which ‘atrophises’ and ‘dissolves’ in the background (pun intended). The ADG-field dynamics is purely algebraic, smooth geometrical base spacetime manifoldless, connection field-solipsistic and autonomous (:dynamical connection field ‘self-governing’ and ‘self-propagating’). [^58] In this respect, it has been noted and emphasised elsewhere [55, 56, 70, 72, 73] that the ‘unitary’ ADG-theoretic field-pairs $\mathcal{F}=(\mathbf{\mathcal{E}},\mathcal{D})$ recall and conceptually resemble Leibniz’s Monads [34], in the sense that they are autonomous (:self-governing), dynamically self-propagating and self-sustaining entities in no need of an external geometrical spacetime continuum for their dynamical support and substenance. Moreover, in close analogy to Leibniz’s purely algebraic (:relational) conception of the notion of derivative in Differential Calculus [13, 7], the ADG-theoretic connection fields are the sources of differentiation, hence the ‘causes’ of dynamical variability in Mallios’s theory. Read below for more detailed discussion of the close resemblance between Mallios’s homological-algebraic conception of Differential Geometry in the guise of ADG and Leibniz’s relational conception of derivative and, in extenso, of Differential Calculus.
We close this subsection by recalling from Brian Hatfield’s prologue to Feynman’s Lectures of Gravitation [15] where he discusses Feynman’s prophetic intuition, vis-à-vis Quantum Gravity, that, in a strong sense, it was quite accidental that gravity was originally formulated in terms of a metric tensor—so that Quantum Gravity would have to involve some kind of ‘quantising spacetime geometry’—but rather that gravity fundamentally reflects some kind of deep gauge invariance. In other words, that gravity should be regarded as a gauge theory like the other three fundamental forces of Nature:
…Thus it is no surprise that Feynman would recreate general relativity from a non-geometrical viewpoint. The practical side of this approach is that one does not have to learn some ‘fancy-schmanzy’ (as he liked to call it) differential geometry in order to study gravitational physics. (Instead, one would just have to learn some quantum field theory.) However, when the ultimate goal is to quantize gravity, Feynman felt that the geometrical interpretation just stood in the way. From the field theoretic viewpoint, one could avoid actually defining—up front—the physical meaning of quantum geometry, fluctuating topology, space-time foam, etc., and instead look for the geometrical meaning after quantization…Feynman certainly felt that the geometrical interpretation is marvellous, ‘but the fact that a massless spin-$2$ field can be interpreted as a metric was simply a coincidence that might be understood as representing some kind of gauge invariance’ [^59] Our emphasis of Feynman’s words as quoted by Hatfield.
2.1.1 The Functoriality of 3rd Quantization in ADG-GT
In ADG, the functoriality that pervades both the kinematical structures and the the purely gauge theoretic dynamical Einstein and Yang-Mills field equations as we expounded above, also extends naturally to include geometric (pre)quantization and second (field) quantization, processes which have also been developed entirely homologically-algebraically by ADG-theoretic means [40, 45, 44, 48, 51, 50, 72, 73]. The upshot of that homological-algebraic quantisation ‘procedure’ is that:
Aphorism 3: The ADG-Theoretic Functorial Quantum Field-Particle Duality. The ADG-theoretic field pair $(\mathbf{\mathcal{E}},\mathcal{D})$ also embodies the fundamental ‘Quantum Field-Particle Duality’ in the sense that from a geometric (pre)quantization and second quantisation vantage, the (local) sections of the vector sheaves $\mathbf{\mathcal{E}}$ embodied in the ADG-theoretic connection fields represent quantum states of bosonic or fermionic field-quanta, as follows: [^60] The correspondences below are borrowed almost verbatim from [40, 45, 44, 72, 73].
$\begin{array}[]{c}(\mathcal{D},\mathbf{\mathcal{E}})\Longleftrightarrow(\mathrm{`dynamics^{\prime}},\mathrm{`kinematics^{\prime}})\cr(\mathcal{D},\Gamma(U,\mathbf{\mathcal{E}}))\Longleftrightarrow(\mathrm{field},\mathrm{local\,particle\,states})\cr(\mathcal{D},\Gamma(U,\mathcal{L}_{n=1})\Longleftrightarrow(\mathrm{Boson\,field},\mathrm{local\,bosonic\,particle\,states})\cr(\mathcal{D},\Gamma(U,\mathbf{\mathcal{E}}_{n>1})\Longleftrightarrow(\mathrm{Fermion\,field},\mathrm{local\,fermionic\,particle\,states})\cr\end{array}$ (32)
Moreover, the correspondences above have been seen to be purely functorial [40, 45, 44, 72, 73]. [^61] Additionally, the reader should note that the fermionic sheaves $\mathbf{\mathcal{E}}_{n>1}$ in the last line above may be conventionally regarded as ‘Odd-Grassmannian’ Sheaves [18, 23].
2.2 Demistifying and Circumventing Singularities and Field-Theoretic Unphysical Infinities.
The $\mathbf{A}$-functorial and $\mathbf{A}$-invariant bottom-up aufbau of ADG has been used to totally circumvent, ‘deconstruct’ and ‘demistify’ singularities and other associated (non-renormalisable) unphysical infinities that have hitherto seemed to mar and assail the CDG-based Einsteinian gravity and the quantum gauge field theories of matter both of which are based on an underlying smooth geometrical spacetime manifold.
The usual physicists’ consensus is that the singularities of General Relativity (GR) and the unphysical infinities of the spacetime continuum based quantum field theories of matter are indications that:
The Laws of Physics break down, thus Nature becomes nonsensical and unpredictable, at those sites [26, 27, 28, 29, 21, 22, 89].
With Professor Mallios we had time and again scrutinised the statement above and invariably it seemed fundamentally incomprehensible to him:
How come the Field Laws of Physics, which we normally model after differential equations, break down at ‘geometrical sites’ we call singularities?
It was Mallios’s original and fundamental idea that:
If we could somehow abstract, genaralise and detach the ‘innate differential geometric mechanism’ of Calculus from its apparent inextricable dependence on a fixed $\mathcal{C}^{\infty}$-smooth background geometrical (spacetime) manifold, we could still do most (if not all!) of Differential Geometry without getting stuck on or breaking down at singularities or other unphysical continuous field infinities. Moreover, we could even integrate, encompass or even ‘absorb’ singularities into the structure sheaf $\mathbf{A}$ of ADG and still the whole differential geometric mechanism would hold in their very presence. That is, the $\mathbf{A}$-connection fields would still define and obey differential equations (laws) in the very presence of singularities, no matter how numerous, robust or pathological those singularities might be, especially when viewed from the vantage of the Classical Differential Calculus (CDG), which is based on a background smooth geometrical spacetime manifold [41, 42, 46, 47, 56, 69, 57, 58].
In other words, [^62] And this was by far Tasos’s favourite motto.
Don’t ‘blame’ Nature for the singularities and the related unphysical infinities of the CDG-based field theories we have. Blame our Mathematics: pit it on CDG itself! [^63] Here, Tasos liked to use the following analogy: in much the same way as the real number line $\mathbb{R}$ comes short or ‘breaks down’ when we try to solve the algebraic equation $x^{2}+1=0$, the CDG-formulated differential equations modelling the Laws of Physics appear to break down at singularities and the latter (misleadingly) appear to be shortcomings and blemishes of the Physical Laws (:differential equations) themselves. However, all we had to do is to extend $\mathbb{R}$ to $\mathbb{C}$, and the ‘seemingly problematic’ $x^{2}+1=0$ is solved(!) Mutatis mutandis then for the extension, abstraction and generalisation of CDG to ADG: from an ADG-theoretic perspective, the Field Laws of Physics do not break down at singularities (see displayed statement below).
Thus, by showing ADG-theoretic means that the differential geometric mechanism of Calculus is inherently or innately homological algebraic (:sheaf and category-theoretic) [^64] More in line with Leibniz’s (rather than Newton’s) conception of the basic derivative operator (viz. connection) of Differential Calculus, Mallios preferred to call ADG a relational theory [13, 7], where the differential geometric mechanism derives from the algebraically modelled (dynamical) relations between the dynamical fields (connections) themselves, not from an external background geometrical spacetime continuum. Read on. and not at all dependent on an external (to the fields) background geometrical spacetime continuum, Mallios abstracts and generalises the usual CDG to ADG and manifestly shows that:
We can actual do/use Differential Calculus in the very presence of singularities and other ‘CDG/differential manifold based anomalies’, thus, a fortiori, the Laws of Physics (:the differential equations between the dynamical connection fields) do not ‘break down’ in any sense at their presence. They still hold intact, and we can still calculate and ‘predict’ things based on them!
We close this subsection by borrowing three quotes from [56]—all three all-time favourites of Tasos—that in a sense foreshadow the development of ADG and its physical applications to gravity and gauge theories of matter and, in view of our arguments in this paper, they ‘post-anticipate’ how our purely algebraic (:ADG-theoretic) finitary, causal and quantal theoresis of vacuum Einstein Gravity and free Yang-Mills theory has come to ‘vindicate’ them:
$\bullet$ The first two quotes by Albert Einstein are taken from the very last Appendix D of The Meaning of Relativity [14]:
…One can give good reasons why reality cannot at all be represented by a continuous field. From the quantum phenomena it appears to follow with certainty that a finite system of finite energy can be completely described by a finite set of numbers. [^65] Our emphasis. This does not seem to be in accordance with a continuum theory, and must lead to an attempt to find a purely algebraic theory for the description of reality [^66] Our emphasis.
and:
…Is it conceivable that a field theory permits one to understand the atomistic and quantum structure of reality? Almost everybody will answer this question with ‘no’. But I believe that at the present time nobody knows anything reliable about it. This is so because we cannot judge in what manner and how strongly the exclusion of singularities reduces the manifold of solutions. We do not possess any method at all to derive systematically solutions that are free of singularities [^67] Our emphasis.
$\bullet$ The third quote by David Finkelstein is taken from the introduction of his Theory of Vacuum [17]:
…The locality principle seems to catch something fundamental about nature… Having learned that the world need not be Euclidean in the large, the next tenable position is that it must at least be Euclidean in the small, a manifold. [^68] Recall that, by definition, a manifold is a locally Euclidean space—that is to say, a space that is locally isomorphic to $\mathbb{R}^{n}$. The idea of infinitesimal locality presupposes that the world is a manifold. [^69] Here, the notion of infinitesimal locality mandates that the Laws of Physics be modelled after differential equations. But the infinities of the manifold (the number of events per unit volume, for example) give rise to the terrible infinities of classical field theory and to the weaker but still pestilential ones of quantum field theory. [^70] Our emphasis. The manifold postulate freezes local topological degrees of freedom which are numerous enough to account for all the degrees of freedom we actually observe.
The next bridgehead is a dynamical topology, in which even the local topological structure is not constant but variable. [^71] Our emphasis. The problem of enumerating all topologies of infinitely many points is so absurdly unmanageable and unphysical that dynamical topology virtually forces us to a more atomistic conception of causality and space-time than the continuous manifold [^72] Again, our emphasis.
2.3 ADG-Field Realism, Solipsism and Monism, and Mallios’s Novel Conception of Bohr’s Correspondence Principle: ADG-Field ‘Unitarity’ from 3rd Quantisation
In the previous section, we witnessed and we argued how the close interplay between fuctoriality and $\mathbf{A}$-invariance of the ADG-field dynamics is tantamount to gauge invariance. That is to say, our attempts to localise and coordinatise (:‘measure’) the Einstein or the Yang-Mills gauge field by employing the algebra structure sheaf $\mathbf{A}$ of generalised coordinates—or ‘arithmetics’, as Mallios preferred to call it—relative to, and localised over, a system of local gauges (:covers) $\mathcal{U}_{i}$ (of the in principle arbitrary base topological space $X$), does not affect the dynamical law that (the curvature of) the field—the $\mathcal{G}\mathcal{M}$-geometric morph of the field—obeys. [^73] That is to say, the action functional density and the law—the differential equation derived from it by ‘variation’ with respect to the local gauge connection potentials $\mathcal{A}$—remains invariant under $\mathbf{A}$-coordinate changes and ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}|_{(U\in\mathcal{U})}$ local gauge transformations.
In turn, this means that that the field ‘sees through’ and remains undisturbed by our ‘perturbing’ generalised acts of measurement (:gauge localisations), hence the issue arises of what would Bohr’s Correspondence Principle of the usual Quantum Theory be in our ADG-GT. Mallios had quite a fascinating, unconventional conception of, and unorthodox ideas about, that, as follows:
Traditionally, from the very advent of Quantum Mechanics, Bohr’s Correspondence Principle can be expressed as follows: observable quantum actions are represented by noncommutative ‘numbers’ (:so-called q-numbers [^74] Think, for instance, of Heisenberg’s matrices.), while our measurements thereof should correspond to (:should yield) commutative numbers (:so-called c-numbers—presumably, these are real numbers in $\mathbb{R}$, which are embedded in the complex numbers $\mathbb{C}$ that the usual Quantum Theory employs [^75] In that sense, real numbers are ‘real’ (pun intended), but over the years, the use of the real number continuum has been questioned and challenged in both Quantum Theory and Quantum Gravity (see for example [30] for a thoughtful exposition).).
Now, Mallios contended that Bohr’s c/q, commutative/noncommutative, classical/quantum dichotomy–the ‘classical/quantum divide’, so to speak—is already embodied and structurally encoded in the ADG-conception of a field as the pair $(\mathbf{\mathcal{E}},\mathcal{D})$, in the following sense:
Our generalised local measurements (:local ‘arithmetics’ relative to a local gauge $U\in\mathcal{U}_{i}$) of the ADG-fields are represented by the local sections of the abelian algebra structure sheaf $\mathbf{A}$ (i.e., $\mathbf{A}_{U}=\mathbf{A}(U)=\Gamma(U,\mathbf{A})$). [^76] The reader should note here a key difference between the usual conception of $c$-numbers (:results of measurements) in conventional Quantum Theory and the generalised $c$-numbers of ADG-GT. The former are sections of the constant sheaf $\mathbf{C}\equiv\mathbb{R}\subset\mathbb{C}$, while the latter are sections of the ‘dynamically variable’ sheaf $\mathbf{A}$, which in turn includes anyway the constant sheaves of complex numbers ($\mathbb{C}$) and real ($\mathbb{R}$) numbers as proper subsheaves. Mallios and Zafiris in [61] give a very novel operational interpretation and physical explanation of this generalisation! On the other hand, the noncommutativity—the $q$-number Heisenberg type of indeterminacy, the quantum fuzziness and the ‘quantum foam’ aspect of the ADG-fields, so to speak—is already encoded in the principal group sheaf ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ of local (gauge) automorphisms of the field, which is locally isomorphic to: ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}(U)\simeq M^{\bullet}_{n}(\mathbf{A}(U))$, the structure group sheaf of invertible $(n\times n)$-matrices, having for entries local sections of $\mathbf{A}$ in $\mathbf{A}|_{U\in\mathcal{U}}=\mathbf{A}(U)\equiv\Gamma(U,\mathbf{A})$. Thus, $M^{\bullet}_{n}(\mathbf{A}(U)$ is the ADG-field theoretic version of the ‘local Heisenberg group’ of the theory, which is manifestly non-abelian—elements (:local sections) of which correspond to the ADG-version of $q$-numbers.
Below, we are going to give briefly a ‘heuristic analogy’ of Mallios’s seemingly unorthodox and unconventional intuition above, which has been previously noted in [55, 72, 73, 56], while in the next subsection we are going to tie our heuristics below to the canonical, sheaf cohomological 3rd quantisation of ADG-GT first proposed in [72].
Similar to how Bohr’s Correspondence is almost tautosemous with the formal ‘inverse’ of the original First Quantization Correspondence:
$\begin{array}[]{c}\mathrm{Classical\,Position\,c-number:}\;x\longrightarrow\hat{x}\,\mathrm{:Quantum\,Position\,q-number}\cr\cr\cr\mathrm{Classical\,Momentum\,c-number:}\;p\longrightarrow\hat{p}\,\mathrm{:Quantum\,Momentum\,q-number}\end{array}$ (33)
also by imposing the following Heisenberg Uncertainty commutation relations between the $\hat{x}$ and $\hat{p}$ operators (:matrices), which in turn generate the usual Heisenberg Algebra of traditional Quantum (Matrix) Mechanics:
$[\hat{x},\hat{p}]=-\imath\hbar$ (34)
the analogous ‘position-momentum correspondence’ within the ADG-fields is, following [55, 72, 73, 56]:
$\begin{array}[]{c}\mathrm{Local\,Particle\,`Position^{\prime}\,States}\longrightarrow\mathrm{Local\,Sections\,of}\,\mathbf{\mathcal{E}}\cr\cr\mathrm{Local\,Field\,`Momentum^{\prime}\,Operator}\longrightarrow\mathcal{D}=d+\mathcal{A}\end{array}$ (35)
the rationale for the heuristic semantic correspondences above being that:
Much in the same way that in conventional particle (Newtonian) mechanics velocity (speed) or momentum is (or measures) the (rate of) change of the position (state) of a particle, [^77] Which momentum is, in turn, the differential (:derivative) of that position determination: $p:\,x\longrightarrow dx$. in ADG-field theory, the $\mathbf{A}$-connection $\mathcal{D}$ in the ADG-field $(\mathbf{\mathcal{E}},\mathcal{D})$ is a generalised differential, acting as a sheaf morphism on the ‘local particle states’ of the ADG-field (which are in turn represented by the local sections of $\mathbf{\mathcal{E}}$ within the ADG-field) as it were to change them.
Then, the formal ‘quantum deformation’ (‘Heisenberg uncertainty relation’) of the usual Poisson Brackets Algebra of Classical Mechanics to the Heisenberg Algebra of Quantum Mechanics by the imposition of the following canonical commutation relations:
$\{x,p\}=0\,\longrightarrow\,[\hat{x},\hat{p}]=-\imath\hbar$ (36)
can heuristically be cast within the ADG-field as follows:
$[\mathbf{\mathcal{E}},\mathcal{D}]={\mathcal{A}}ut\mathbf{\mathcal{E}}\,\stackrel{{\scriptstyle\mathrm{locally}}}{{\longrightarrow}}\,[s\in\mathbf{\mathcal{E}}(U),d+\mathcal{A}]=M_{n}(\mathbf{A}_{U})s=M_{n}(\mathbf{A}(U))s$ (37)
where $s\in\mathbf{\mathcal{E}}_{\mathbf{A}}(U)$ (with $\mathbf{\mathcal{E}}_{\mathbf{A}}(U)\equiv\Gamma(U,\mathbf{\mathcal{E}}_{\mathbf{A}})\simeq\mathbf{A}^{n}(U)$) a ‘local quantum particle state’ within the ADG-field $(\mathbf{\mathcal{E}},\mathcal{D})$ and $M_{n}(\mathbf{A}(U))$ a local section of the structure group sheaf $\mathcal{G}={\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ of the ADG-field relative to the local open gauge $U$.
The ‘canonical’ commutator ‘auto-uncertainty’ relation within the ADG-field $(\mathbf{\mathcal{E}},\mathcal{D})$ in equation (37) above can then be heuristically interpreted as follows:
The ‘canonical’ commutator quantum uncertain relation between the generalised local position quantum particle states in $\mathbf{\mathcal{E}}$ and its dua; generalised momentum field operator (:sheaf morphism) $\mathcal{D}$ generates and induces a dynamical local gauge transformation in ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}|_{U}=M^{n}(\mathbf{A}(U))$, which then acts on the local quantum particle states to change them and, as it were, to ‘blur’ them (:‘quantum fuzziness’ or ‘quantum foam’).    In this heuristic sense, $\mathbf{\mathcal{E}}$ is ‘complementary’ to $\mathcal{D}$, thus the ADG-field may be thought of as being ‘self-complementary’ and ‘self-quantum’. In this sense we argued in [55, 72, 73, 56] that the ADG-field is an already self-quantum, auto-dynamical entity.
We can formalise the generalised ADG-theoretic quantum uncertainty/complementarity relation above as a homological $\mathbf{A}$-tensor product morphism type of map:
$\mathcal{Q}:\,\mathbf{\mathcal{E}}_{\mathbf{A}}\otimes_{\mathbf{A}}\mathcal{D}\longrightarrow{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}_{\mathbf{A}}$ (38)
thus finally arrive at the definition of the ‘unitary’ quantal ADG-gauge field [^78] The reader should note that, issuing from the 3rd Quantisation scheme for ADG-field theory originally presented in [72], the epithets unitary quantum carry standard meaning in the usual Quantum Theory, hence we use inverted single quotes around the word unitary, while instead of quantum we use quantal throughout our work, in order to avoid confusion of semantic reference. as being the following tetrad:
$\mathbf{U}:=(\mathbf{\mathcal{E}},\mathcal{D},{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}},\mathcal{Q})$ (39)
which encodes the following four pieces of important information:
  1. 1.
    The vector sheaf $\mathbf{\mathcal{E}}$ of generalised quantum particle ‘position’ states;
  2. 2.
    The connection field $\mathcal{D}$, which effectuates dynamical ‘momentum’ field-like changes of the said states by acting as a sheaf morphism on $\mathbf{\mathcal{E}}$’s local sections;
  3. 3.
    The structure gauge group sheaf ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ of local gauge transformations, also effectively acting on $\mathbf{\mathcal{E}}$ as its principal structure group sheaf [^79] And conversely, $\mathbf{\mathcal{E}}$ can be viewed as the associated sheaf to the principal group sheaf ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$.; and finally,
  4. 4.
    The quantum uncertainty operator (:morphism) $\mathcal{Q}$, which stands for an ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$-valued quantum act of perturbation on the dynamical action of the connection field on the local quantum particle states (:local sections) of $\mathbf{\mathcal{E}}$.
The epithet ‘unitary’ for $\mathbf{U}$ in (39) above indicates that it is a holistic entity, an inseparable whole, encoding state ($\mathbf{\mathcal{E}}$), dynamical changes of state ($\mathcal{D}$), gauge symmetries and invariances of dynamical changes of state (${\mathcal{A}}ut\mathbf{\mathcal{E}}$), and quantum uncertainty of ‘determination’ ($\mathcal{Q}$), all-4-in-1.
All there is in ADG-field theory are the Leibnizian Monad type of entities $\mathbf{U}=(\mathbf{\mathcal{E}},\mathcal{D},{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}},\mathcal{Q})$, nothing else. This is what was referred to in [55, 72, 73, 56] as ‘ADG-field solipsism and monism’.
Thus, perhaps more importantly, the adjective ‘unitary’ given to the tetrad $\mathbf{U}=(\mathbf{\mathcal{E}},\mathcal{D},{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}},\mathcal{Q})$ above pertains to the fact that in ADG-field theory, there is no reference and recourse to, no dependence whatsoever on, an external (to the ADG-fields themselves) background geometrical spacetime manifold. It follows that the usual distinction and schism internal/external that is normally reserved for the symmetries of the usual gauge field theories of matter—whether classical or quantum—loses its meaning in our ADG-theoresis of Vacuum Einstein Gravity and Free Yang-Mills theories. [^80] Traditionally, we reserve the epithet ‘external’ for ‘the external (to the fields) spacetime symmetries’, while ‘internal’ is normally reserved for gauge degrees of freedom and their symmetries [4, 23]. This has also been succinctly pointed out more recently in [61].
2.3.1 Interregnum: Drawing Formal Links with 3rd Quantisation
In the discussion below, we briefly draw links between the formal ‘canonical’ quantisation heuristics above and the functorial sheaf cohomological 3rd Quantisation of ADG-GT scenario presented in [72].
To that end, we recall that in [72] we intuited that since the ADG-fields $\mathcal{F}=(\mathbf{\mathcal{E}},\mathcal{D})$ are dynamically self-supporting, autonomous monadic entities as we emphasised earlier, and moreover, since they are ‘self-dual [^81] In the sense that the connection momentum-like field $\mathcal{D}$ is quantum dual to the ‘position quantum particle states’ (represented by the local sections of) $\mathbf{\mathcal{E}}$.
a possible quantization scenario for them should involve solely their two constitutive parts, namely, $\mathbf{\mathcal{E}}$ and $\mathcal{D}$, without recourse to/dependence on extraneous structures (e.g., a base spacetime manifold) for its mathematical support and its self-consistent (physical) interpretation.
Thus, in what formally looked like a canonical quantization-type of scenario,
in [72] we envisaged abstract non-trivial local commutation relations between the abstract position ($\mathbf{\mathcal{E}}$) and momentum (:$\mathcal{D}$) aspects of the ADG-fields.
To that end, we recalled that
there are certain local (:differential) forms that uniquely characterize sheaf cohomologically the vector sheaf $\mathbf{\mathcal{E}}$ and the connection $\mathcal{D}$ parts of the ADG-fields $\mathcal{F}=(\mathbf{\mathcal{E}},\mathcal{D})$
Thus, the basic heuristic-intuitive idea in [72] was to identify the relevant forms and then posit non-trivial commutation relations between them. Moreover, for the sake of the aforementioned ‘dynamical ADG-field autonomy’, we would like to require that
the envisaged commutation relations should not only involve just the two components (i.e., $\mathbf{\mathcal{E}}$ and $\mathcal{D}$) of the total ADG-fields $\mathcal{F}=(\mathbf{\mathcal{E}},\mathcal{D})$, but they should also somehow ‘algebraically close’ within the fields themselvesi.e., the result of their commutation relations should not take us ‘outside’ the total ADG-field structure (and its ‘dynamical auto-transmutations’), which anyway is the only dynamical structure involved in our theory. [^82] This loosely reminds one of the theoretical requirement for algebraic closure of the algebra of quantum observables in canonical QG, with the important difference however that the $\mathrm{Diff}(M)$ group of the external (to the gravitational field) spacetime manifold must also be considered in the constraints, something that makes the desired closure of the observables’ algebra quite a hard problem to overcome [82]. In [55, 72, 73] we discuss certain difficult problems that $\mathrm{Diff}(M)$ creates in various QG approaches, as well as how its manifest absence in ADG-gravity can help us bypass them totally. For, recall that from the ADG-perspective gravity is an external (:background) spacetime manifold unconstrained (because it is a background spacetime manifoldless) pure gauge theory (:of the 3rd kind).
Keeping the theoretical requirements above in mind, we recall from [38, 39, 40, 54, 44] two important sheaf cohomological results:
  1. 1.
    That, sheaf cohomologically, the vector sheaves $\mathbf{\mathcal{E}}$ are completely characterized by a so-called coordinate $1$-cocycle $\phi_{\alpha\beta}\in Z^{1}(\mathcal{U},{\mathcal{G}}{\mathcal{L}}(n,\mathbf{A}))$ associated with any system $\mathcal{U}$ of local gauges of $\mathbf{\mathcal{E}}$. Intuitively, this can be interpreted in the following ‘Kleinian symmetry-geometry’ way: since any (vector) sheaf is completely determined by its (local) sections, [^83] A basic motto (:fact) in sheaf theory is that “a sheaf is its sections[6, 38]. If we know the local data (:sections), we can produce the whole sheaf space by restricting and collating them relative to an open cover $\mathcal{U}$ of the base topological space $X$. This is the very process of ‘sheafification’ (of a preasheaf) [6, 38]. one way of knowing the latter is to know how they transform—in passing, for example, from one local gauge ($U_{\alpha}\in\mathcal{U}$) to another ($U_{\beta}\in\mathcal{U}$), with $U_{\alpha}\cap U_{\beta}\neq\emptyset$ and $\mathcal{U}$ a chosen system of local open gauges covering $X$ [^84] In particular, $\phi_{\alpha\beta}$ can be locally expressed as the $\mathbf{A}|_{U_{\alpha\beta}}$-isomorphism: $\phi_{\alpha}\circ\phi_{\beta}^{-1}\in{\mathcal{A}}ut_{\mathbf{A}_{\alpha\beta}}(\mathbf{A}^{n}|_{U_{\alpha\beta}})=\mathrm{GL}(n,\mathbf{A}(U_{\alpha\beta}))={\mathcal{G}}{\mathcal{L}}(n,\mathbf{A})(U_{\alpha\beta})$, in which expression the familiar local coordinate transition (:structure) functions appear. Hence, also the ‘natural’ structure (:gauge) group sheaf ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}={\mathcal{G}}{\mathcal{L}}(n,\mathbf{A})$ of $\mathbf{\mathcal{E}}$ arises. To know something (e.g., a ‘space’) is to know how it transforms, the fundamental idea underlying Klein’s general conception of ‘geometry’ [32].
    Thus, the bottom-line here is that the characteristic cohomology classes of vector sheaves $\mathbf{\mathcal{E}}$ are completely determined by $\phi_{\alpha\beta}$; write:
    $[\phi_{\alpha\beta}]\in H^{1}(X,{\mathcal{G}}{\mathcal{L}}(n,\mathbf{A}))=\lim_{\overrightarrow{\;\;\mathcal{U}\;\;}}H^{1}(\mathcal{U},{\mathcal{G}}{\mathcal{L}}(n,\mathbf{A}))$ (40)
    where the $\mathcal{U}$s, normally assumed to be locally finite open coverings of $X$ [38, 39, 53, 54, 55], constitute a cofinal subset of the set of all proper open covers of $X$ [^85] An assumption that has proven to be very fruitful in applying ADG to the formulation of a locally finite, causal and quantal Vacuum Einstein Gravity and Free Yang-Mills theories, as we argued in the first part of this paper and throughout our past works [53, 54, 55, 69, 70, 71, 72, 73]. En passant, we also note that in [72] the direct (:inductive) limit depicted in (40) above is secured by the ‘cofinality’ of the set of finitary (:locally finite) open coverings of $X$ that we choose to employ [78, 74, 75, 67, 68, 53, 54, 55, 69, 70] and it was emloyed $K$-theoretically to link the 3rd Quantisation of ADG-fields scenario with Mallios’s $K$-theoretic musings on topological algebra structure sheaves $\mathbf{A}$ and the 2nd Quantisation classification of the local quantum particle states (:local sections) of the vector sheaves involved in the ADG-fields into Bosons ($\mathbf{\mathcal{E}}$ is a line sheaf $\mathcal{L}_{n=1}$) and Fermions ($\mathbf{\mathcal{E}}$ is a vector sheaf of rank $n>1$) [45]. In toto, we assume that $\phi_{\alpha\beta}$ encodes all the (local) information we need to determine the local quantum-particle states of the field in focus (i.e., the local sections of $\mathbf{\mathcal{E}}$).
  2. 2.
    On the other hand, it was observed in [72] that locally $\mathcal{D}$ is uniquely determined by the so-called ‘gauge potential’ $\mathcal{A}$, which is normally (i.e., in CDG) defined as a Lie algebra (:vector) valued $1$-form [23]. Correspondingly, in ADG $\mathcal{A}$ is seen to be an element of $M_{n}(\Omega(U))=M_{n}(\Omega)(U)=\Omega(\mathbf{\mathcal{E}}nd\mathbf{\mathcal{E}})$, [^86] Note that, as also mentioned earlier, in ADG by definition one has: $\Omega:=\mathbf{\mathcal{E}}^{*}:=\mathcal{H}om_{\mathbf{A}}(\mathbf{\mathcal{E}},\mathbf{A})$. That is, the $\mathbf{A}$-module sheaf $\Omega$ of abstract differential $1$-forms is dual to the vector sheaf $\mathbf{\mathcal{E}}$, much like in the classical theory (:CDG of $\mathcal{C}^{\infty}$-smooth manifolds) where differential forms (:cotangent vectors) are dual to tangent vectors [23], although as it has been empasised throughout our works, in ADG the epithet ‘(co)tangent’ is meaningless due to the manifest absence of an operative background space(time) of any kind (and especially of a base manifold). thus it is called the local $\mathbf{A}$-connection matrix $(\mathcal{A}_{ij})$ of $\mathcal{D}$, with entries local sections of *$\mathbf{\mathcal{E}}^{*}=\Omega$. In turn, this means that locally $\mathcal{D}$ splits in the familiar way, as follows:
    $\mathcal{D}=d+\mathcal{A}$ (41)
    where $\partial$ is the usual ‘inertial’ (:flat) differential [^87] As noted in the previous section, in ADG, the Cartan-Kähler differential $d$, like $\mathcal{D}$, is defined as a linear, Leibnizian $\mathbf{C}$-sheaf morphism $d:\;\mathbf{A}\rightarrow\Omega$, thus it is an instance of on $\mathbf{A}$-connection; albeit, a flat one (:$R(d)=d^{2}=0$), which is secured by the very definition of curvature in (17) and the niloptency of the exterior differential $d$. and $\mathcal{A}$ the said gauge potential. In ADG-gravity, the total field $\mathcal{D}$ as a whole (:‘globally’) represents the gravito-inertial field, but locally it can be separated into its inertial (:$d$) and gravitational (:$\mathcal{A}$) parts. [^88] In the classical theory of gravity (:General Relativity; abbr. GR), the physical meaning of this local separation of the total field $\mathcal{D}$ into $\partial$ and $\mathcal{A}$ reflects the local principle of equivalence; namely, that locally, the spacetime manifold $M$ of GR is flat Minkowski space, or equivalently, that locally, GR reduces to Special Relativity, or perhaps more importantly, that gravity can always be ‘gauged away’ locally by a suitable choice of ‘gauge’ (:local inertial frame). This is simply Einstein’s elevator gedanken experiment [63].
Thence, the envisaged sheaf cohomological canonical quantization-type of scenario for the total ADG-fields $\mathcal{F}=(\mathbf{\mathcal{E}},\mathcal{D})$ rests essentially on positing the following non-trivial abstract Heisenberg-type local commutation relations between (the characteristic forms that completely characterize) $\mathbf{\mathcal{E}}$ (:abstract ‘position’ particle states) and $\mathcal{D}$ (:abstract ‘momentum’ field operator). Thus, heuristically we posited in [72] the following ‘canonical’ commutation relations:
$[\phi,\mathcal{D}]\stackrel{{\scriptstyle loc.}}{{=}}[\phi_{\alpha\beta},d+\mathcal{A}_{ij}]_{U_{\alpha\beta}}=[\phi_{\alpha\beta},d]_{U_{\alpha\beta}}+[\phi_{\alpha\beta},\mathcal{A}_{ij}]_{U_{\alpha\beta}}$ (42)
stressing also that, as highlighted in [72],
the local commutation relations in (42) above are well defined, since they effectively close within the noncommutative $(n\times n)$-matrix Klein-Heisenberg algebra $\mathbf{\mathcal{E}}nd\mathbf{\mathcal{E}}(U_{\alpha\beta})=M_{n}(\mathbf{A}(U_{\alpha\beta}))=M_{n}(\mathbf{A})(U_{\alpha\beta})$ of the field’s endomorphisms—the field’s ‘noncommutative Kleinian geometry’ we mentioned earlier representing what Mallios intuited as some kind of ‘quantum field foam’—the intrinsically noncommutative aspect of the ADG-fields.
This ‘algebraic closure’ is in accord with the theoretical requirement we imposed earlier, namely that,
the abstract, Heisenberg-like, canonical quantum commutation relations between the two components $\mathbf{\mathcal{E}}$ and $\mathcal{D}$ of the ADG-fields should not take us outside the fields, but should rather close within them. [^89] Here, one could envisage an abstract Heisenberg-type of algebra freely generated (locally) by $\phi$ (:abstract position) and $\mathcal{A}$ (:abstract momentum), modulo the (local) commutation relations (41). Plainly, it is a subalgebra of $\mathbf{\mathcal{E}}nd\mathbf{\mathcal{E}}(U)$, but deeper structural investigations on it must await a more complete and formal treatment.
Indeed, $\mathbf{\mathcal{E}}nd\mathbf{\mathcal{E}}$ is precisely the algebra sheaf of internal/intrinsic (dynamical) self-transmutations of the (quantum particle states of the) field—by definition, the $\mathbf{\mathcal{E}}$-endomorphisms in ${\mathcal{H}}om_{\mathbf{A}}(\mathbf{\mathcal{E}},\mathbf{\mathcal{E}})$ (:quantum field foam).
This is another aspect of the quantum dynamical autonomy of ADG-fields:
the $\mathbf{\mathcal{E}}$ (:abstract point-particle/position) part of the ADG-field is ‘complementary’, in the quantum sense of ‘complementarity’, to $\mathcal{D}$ (:abstract field-wave/momentum). Thus, the total ADG-fields =(.$\mathcal{F}=(\mathbf{\mathcal{E}}.\mathcal{D})$ are ‘quantum self-dual’ entities [55, 56, 69, 72]. [^90] From our abstract and background spacetime manifoldless perspective, the de Broglie-Schrödinger wave-particle duality is almost tautosemous with the Bohr-Heisenberg momentum-position complementarity.
Furthermore, by choosing $\phi_{ab}=\phi_{ab}^{in}$ [^91] The superscript ‘$in$’ stands for ‘inertial’, and it represents a choice (:our choice!) of a local change-of-gauge $\phi^{in}_{\alpha\beta}\in\mathcal{G}\mathcal{L}(n,\mathbf{A})_{\alpha\beta}\equiv\Gamma(U_{\alpha\beta},\mathcal{G}\mathcal{L}(n,\mathbf{A})$ that would take us to a locally inertial frame of $\mathbf{\mathcal{E}}$ over $U_{\alpha\beta}\subset X$. so that $\mathcal{A}$ is ‘gauged away’—i.e., by setting $\mathcal{A}=0$, [^92] As noted earlier, this is an analogue of the Equivalence Principle (EP) of GR in ADG-gravity, corresponding to the local passage to an ‘inertial frame’ (:one ‘covarying’ with the gravitational field; e.g., recall Einstein’s free falling elevator gedanken experiment) in which the curved gravito-inertial $\mathcal{D}$ reduces to its flat ‘inertial’ $\mathbf{A}$-connection part $d$ [38, 39, 53, 54, 55]. As noted above, this just reflects the well known fact that GR is locally SR, or conversely, that when SR is localized (ie, ‘gauged’ over the base spacetime manifold) it produces GR (equivalently, the curved Lorentzian spacetime manifold of GR is locally the flat Minkowski space of SR). In summa, gravity (:$\mathcal{A}$) has been locally gauged away, and what we are left with is the inertial action $d$ of the ADG-gravitational field $\mathcal{D}$. It must be also stressed here that the choice of a locally inertial frame, like all gauge choices, is an externally imposed constraint in the theory—‘externally’, meaning that it is we, the external (to the field) experimenters/theoreticians (‘observers’) that we impose such constraints on the field (i.e., we make choices about what aspects of the field we would like to single out and, ultimately, observe/study). reduces (41) to (omitting the local open gauge indices/subscripts ‘$\alpha,\beta$’):
$[\phi^{in},d]=\phi^{in}\circ d-d\circ\phi^{in}$ (43)
Moreover, since we are sheaf cohomologically guaranteed that $d\circ\phi=0$ globally, which is tantamount to the very existence of an $\mathbf{A}$-connection $\mathcal{D}$ (globally) on $\mathbf{\mathcal{E}}$ [38, 39, 44], [^93] This essentially corresponds to the fact that the coordinate $1$-cocycle $\phi_{\alpha\beta}\in Z^{1}(\mathcal{U},\Omega)$ is actually a coboundary (:a closed form), belonging to the zero cohomology class $[d\phi_{\alpha\beta}]=0\in H^{1}(X,M_{n}(\Omega))$, which in turn guarantees the existence of an $\mathbf{A}$-connection on $\mathbf{\mathcal{E}}$ as the so-called Atiyah class $\mathbf{A}$ of $\mathbf{\mathcal{E}}$ vanishes (:$\mathbf{A}(\mathbf{\mathcal{E}}):=[d\phi_{\alpha\beta}]=0$) [38, 39, 44]. (42) further reduces to:
$[\phi^{in},d]=\phi^{in}\circ d$ (44)
Now, a heuristic physical interpretation can be given to (43) if we consider its effect (:action) on a local section $s\in\mathbf{\mathcal{E}}_{\alpha\beta}:=\mathbf{\mathcal{E}}(U_{\alpha\beta})\equiv\mathbf{\mathcal{E}}|_{U_{\alpha\beta}}$:
$[\phi^{in},d](s)=(\phi^{in}\circ d)(s)=\phi^{in}(ds)$ (45)
(44) designates the inertial dynamical action of $\mathcal{D}$ (i.e., the action of its locally flat, inertial part $d$) on (an arbitrary) $s$, followed by the gauge transformation of $ds$ to an inertial frame $e_{in}^{U_{\alpha\beta}}\subset\mathbf{\mathcal{E}}_{\alpha\beta}$covarying’ with the inertio-gravitational field.
It may be interpreted as expressing what happens to a ‘vacuum graviton state’ $s$ when it is first acted upon [^94] Recall that we are considering only vacuum gravity, in which the non-linear gravitational field ‘couples’ solely to itself(!) by the inertial part of the total ADG-gravitational field $\mathcal{D}$ and then [^95] The sequential language used here should not be interpreted in an temporal-operational sense—as it were, as ‘operations carried out sequentially in time’. to an inertial frame that in a sense ‘covaries’ with the said inertial change $d$ of $s$.
Heuristically, we further intuited in [72] that one can perhaps get a more adventurous (meta)physical insight into (44) by defining the uncertainty operator $\mathcal{U}$ as
$\mathcal{U}:=\phi^{in}\circ\partial\in\mathbf{\mathcal{E}}nd\mathbf{\mathcal{E}}$ (46)
and by delimiting all the quantum-particle (:abstract position) states of the field (:local sections of $\mathbf{\mathcal{E}}$) that are annihilated by it. Intuitively, these are formally the local ‘classical-inertial’ states
$\mathbf{\mathcal{E}}^{cl}_{U}:=\mathrm{span}_{\mathbf{\mathbb{C}}}\{s\in\mathbf{\mathcal{E}}(U):\;\mathcal{U}(s)=0\}=:\mathrm{ker}(\mathcal{U})$ (47)
for which the abstract sheaf cohomological Heisenberg uncertainty relations (42) vanish. Plainly, $\mathbf{\mathcal{E}}^{cl}(U)$ is a $\mathbf{C}$-linear subspace of $\mathbf{\mathcal{E}}(U)$the kernel of $\mathcal{U}$.
On top of the above, intuitively it makes sense to assume that $\mathcal{U}$ is a ‘projector’—a primitive idempotent (:projection operator) locally in $\mathbf{\mathcal{E}}nd\mathbf{\mathcal{E}}$ (i.e., in $M_{n}(\mathbf{A}(U))$)—since the ‘gedanken’ operation of ‘inertially covarying with a chosen local inertial frame’ must arguably be idempotent. [^96] After all, ‘inertially covarying the inertial state leaves it inertially covariant’. Or, to use a famous Einstein ‘gedanken metaphor’: ‘jumping on a light-ray (in order to ride it) twice, simply leaves you riding it’(!) This means that $\mathcal{U}^{2}=\mathcal{U}$, so that $\mathcal{U}$ separates (chooses or projects out) the ‘classical’ ($\mathrm{eigen}_{0}(\mathcal{U})\equiv\mathrm{ker}(\mathcal{U})$) from the quantum ($\mathrm{eigen}_{1}(\mathcal{U})$) local quantum gravito-inertial states. [^97] In [72], a formal mathematical reason why we chose $\mathcal{U}$ to be a projection operator was to apply it and relate our 3rd Quantisation scenario to Mallios’s $K$-theoretic perspective on 2nd (Field) Quantisation in [45]..
Finally, in line with [72], we would like to ask en passant here the following highly speculative question:
Could the generation/emergence of (inertio-gravitational) mass be somehow accounted for by a (spontaneous) symmetry breaking-type of mechanism, whereby, the dynamical automorphism group ${\mathcal{A}}ut\mathbf{\mathcal{E}}$ of the ADG-gravitational field $(\mathbf{\mathcal{E}},\mathcal{D})$ reduces to its subroup that leaves $\mathrm{ker}(\mathcal{U})$ invariant? Alternatively intuited, could the emergence of inertio-gravitational mass be thought of as the result of some kind of ‘quantum anomaly’ of 3rd-quantized Vacuum Einstein Gravity? [^98] The epithet ‘quantum’ adjoined to ‘anomaly’ is intended to distinguish the effect intuited above from the usual anomalies. A ‘quantum anomaly’ is the ‘converse’ of an anomaly in the usual sense, in that what was a symmetry of the quantum theory (:an element of ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ in our case) ceases to be a symmetry of the ‘classical domain’ of our theory (:$\mathrm{ker}(\mathcal{U})$). Let it be stressed that the emergence of gravito-inertial ‘mass’ in the sense intuited here has a truly relational (:algebraic) and ‘global’ flavour reminiscent of Mach’s ideas: ‘global’ gravitational field symmetries in ${\mathcal{A}}ut\mathbf{\mathcal{E}}$ are locally reduced to inertial ones, and sheaf theory’s ability to interplay between local and global comes in handy in this respect [61]. (In [61], Mallios and Zafiris do a great job in highlighting exactly how sheaf theory allows one to go from ‘local’ to ‘global’, and vice versa.)
$\bullet$ Identifying 3rd Quantisation within our ‘Unitary’ Quantal ADG-Gauge Field Tetrads. The alert reader must have already noticed that the ‘canonical’ commutation relations (42) are the sheaf cohomological versions of our ‘heuristic’ canonical commutation relations (37). Also, by comparing the commutator expressions (37) and (43) and the associated definitions of the operators (:morphisms) $\mathcal{Q}$ in (38) and $\mathcal{U}$ in (46), the astute reader must have realised that $Ran(\mathcal{Q})\subset Ran(\mathcal{U})$, as $M_{n}^{\bullet}(\mathbf{A}(U))\subset\mathbf{\mathcal{E}}nd\mathbf{\mathcal{E}}|_{U}=M_{n}(\mathbf{A}(U))$.
Overall, and without loss of generality of mathematical structures or physical interpretation thereof, in the light of 3rd Quantisation our ‘Unitary’ Quantal ADG-Gauge Field Tetrads in (39) can now be identified with the tetrad:
$\mathbf{U}:=(\mathbf{\mathcal{E}},\mathcal{D},{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}},\mathcal{Q})\equiv(\mathbf{\mathcal{E}},\mathcal{D},\mathbf{\mathcal{E}}nd\mathbf{\mathcal{E}},\mathcal{U})$ (48)
which carries the same denomination as before ‘Unitary’ Quantal ADG-Gauge Field and is again symbolised by $\mathbf{U}$.
Now, by recalling the meaning we ascribed to $\mathbf{U}$ as an inseparable (:‘unitary’), self-dual, 3rd-quantised, auto-dynamical, background spacetime manifoldless gauge field of the 3rd kind, we can, mutatis mutandis address the traditional distinction and schism in the usual quantum theory between ‘observer’ (classical exosystem) and ‘observed’ (quantum endosystem) [90, 17]. [^99] This pertains to the (in)famous Heisenberg scnitt (:Heisenberg cut): the schism that divides and separates the classical from the quantum phenomena and it delimits the boundary across which the wave function supposedly collapses upon measurement. It is when $q$-numbers become $c$-numbers, and probability amplitudes become probability distributions. See [90] for a plethora of classic articles on the Heisenberg Schnitt and the Quantum Theory of Measurement. It’s where John Wheeler noted that “No phenomenon is a phenomenon unless it is an observed phenomenon[90]. We are not going to elaborate in detail on this subtle and important point here,100100100We are going to tackle it further in [56]. but it is noteworthy to mention that significant work has been done on defining by the homological algebraic (:category and topos-theoretic) means of ADG internally consistent quantum observables in the theory, without recourse to any external spacetime manifold, as befits the ‘unitary’ and quantal ADG-gauge field theory [92, 93, 61, 61].
In the light of the physical interpretation of the $\mathbf{U}$ ADG-field tetrads above, we conclude this subsection by quoting David Finkelstein from [17], making some protphetic remarks about the future of physical laws vis-à-vis his Quantum Relativity Theory approach to Quantum Gravity, based on ‘abolishing’ this external/internal field distinction and schism of the usual theory:101101101This quote was one of Tasos’s all-time favourites, and I recall fondly him urging me to include it in the opening talk of the first Glafka 2004: Iconoclastic Approaches to Quantum Gravity international theoretical physics conference that we jointly organised in Athens, Greece—setting thus ‘The Spirit of the Meeting[71].
“…What are we after as physicists? Once I would have said, the laws of nature; then, the law of nature. Now I wonder.102102102Our emphasis.
A law, or to speak more comprehensively, a theory, in the ordinary sense of the word, even a quantum theory of the kind studied today by almost all quantum physicists, is itself not a quantum object. We are supposed to be able to know the theory completely, even if it is a theory about quanta. Its symbols and rules of inference are supposed to be essentially non-quantum. For example, ordinary quantum theory assumes that we can know the form of the equations obeyed by by quantum variables exactly, even though we cannot know all the variables exactly. This is considered consistent with the indeterminacies of quantum theory, because the theory itself is assumed to sum up conclusions from arbitrarily many experiments.
Nevertheless, since we expect that all is quantum, we cannot consistently expect such a theory to exist except as an approximation to a more quantum conception of a theory. At present we have non-quantum theories of quantum entities. Ultimately the theory too must reveal its variable nature. For example, the notion that an experiment can be repeated infinitely often is as implausible as the notion that it can be done infinitely quickly ($c=\infty$), or infinitely gently ($\hbar=0$).
It is common to include in the Hamiltonian of (say) an electron a magnetic field that is treated as a non-quantum constant, expressing the action of electric currents in a coil that is not part of the endosystem but the exosystem. Such fields are called external fields. Upon closer inspection, it is understood, the external field resolves into a host of couplings between the original electron and those in the coil system, now part of the endosystem.
It seems likely that the entire Hamiltonian ultimately has the same status that we already give the external field. No element of it can resist resolution into further quantum variables. In pre-quantum physics the ideal of a final theory is closely connected with that of a final observer, who sees everything and does nothing. The ideal of a final theory seems absurd in a theory that has no final observer. When we renounce the ideal of a theory as a non-quantum object, what remains is a theory that is itself a quantum object. Indeed, from an experimental point of view, the usual equations that define a theory have no meaning by themselves, but only as information-storing elements of a larger system of users, as much part of the human race as our chromosomes, but responding more quickly to the environment. The fully quantum theory lies somewhere within the theorizing activity of the human race itself, or the subspecies of physicists, regarded as a quantum system. If this is indeed a quantum entity, then the goal of knowing it completely is a Cartesian fantasy, and at a certain stage in our development we will cease to be law-seekers and become law-makers.
It is not clear what happens to the concept of a correct theory when we abandon the notion that it is a faithful picture of nature. Presumably, just as the theory is an aspect of our collective life, its truth is an aspect of the quality of our life103103103Again, our emphasis throughout.
3 Brief Philosophical Epilegomena: Anastasios Mallios’s Original Vision and Posthumous Future Legacy
This author’s earliest recollections of exchanges with Professor Anastasios Mallios in the late 1990s/early 2000s during multiple dinner evenings at a little known, quaint and cosy little tavern, fittingly called Algebra, situated in the northern Athens suburb of Paleo Psychiko, about the potential import of Abstract Differential Geometry in current persistently un(re)solved technical (:mathematical), conceptual-cum-semantic and philosophical issues in Quantum Gravity and Quantum Gauge Theory research, focused mainly on two fronts:
  1. 1.
    The algebraic essence and origin of ‘physical space’ and its ‘physical geometry’ and, as a ‘result’:
  2. 2.
    The non-existence of an a priori ‘geometrical space(time)’, but quite on the contrary, the emergence of ‘geometrical space(time)’ as an outcome of an algebraic (:relational) dynamics (:dynamical interactions) between the ‘physical geometrical objects’ (:the physical fields) that live on a surrogate and virtual ‘space’.
Focusing on the two items above, below I will try to recall and ‘reconstruct’ the origins and motivations of ADG.
3.1 The Original Vision: ‘Geometrical Space(-Time)’ comes from ‘Algebraic Dynamics’
From numerous exchanges, close collaboration and warm friendship with the creator of ADG over more than one and a half decades, this author maintains and is willing to argue that Professor Mallios’s inspired magnum opus (ADG) was originally motivated by two main aspects—one technical-mathematical, the other intuitive-heuristic-conceptual and physical—that synergistically feed each other, grow holistically together, and almost perfectly dovetail with each other in the aufbau of ADG..
3.1.1 Tracing the Mathematical ‘Origins’ of ADG: the Categorical Duality between Algebra and Geometry
Very early on in my crossing worldlines with Professor Mallios, he kept on bringing up one of his all-time favourite quotes by Sophie Germain [20] as being a motto at the very heart of ADG:
Géometrie est une Algèbre bien figuré, mais Algèbre est une Géometrie bien écrite104104104English translation: “Geometry is a well figured (or designed) Algebra, while Algebra is a well written Geometry”.
The quote above perfectly encapsulates, in a beautifully poetic way, the fundamental mathematical (:categorical) duality between Algebra and Geometry, which, in modern mathematical (:category-theoretic) parlance may be boiled down and reduced to two cornerstone results, both of which we have played a central role in Mallios’s developing his theory:105105105Tasos Mallios in numerous private communications.
  1. 1.
    Gel’fand Duality and the Gel’fand-Stone Theorem: Generally and loosely speaking, Gel’fand duality is a general duality between spaces and algebras of functions defined on them. In particular, for the case of compact topological spaces and abelian $C^{*}$-algebras, Gel’fand duality roughly pertains to the result that every commutative $C^{*}$-algebra $A$ is equivalent to the abelian $C^{*}$-algebra of continuous functions on a suitably and ‘naturally topologised’ (:using the algebraic structure itself)106106106The set of the algebra’s idecomposable, irreducible atomic elements so to speak—its (primitive or prime) ideals. space called its Gel’fand Spectrum $Spec(A)$ [19, 31, 33].107107107Earlier in the present paper, we witnessed an instance of ‘discrete’ Gel’fand Duality, when we discussed Sorkin’s finitary poset discretisations of locally compact continuous manifolds and their Gel’fand-dual incidence Rota algebras (cf. Section 1). It must be emphasised here that the adjective ‘equivalent’ in the statement of Gel’fand Duality above pertains to the categorical equivalence (:functorial correspondence) between the category of abelian $C^{*}$-algebras and that of (compact) topological spaces.
    For Tasos, the main technical and conceptual essence of the mathematical result above is that ‘Space (Geometry/Topology) can be somehow ‘derived’ or extracted from Algebraic Structure’.108108108Again, Tasos Mallios in numerous private communications. In [43], for example, Tasos goes at great lengths, by using Gel’fand Duality and the so-called Gel’fand Transform of an algebra, in delimiting the sort of (topological) algebra sheaves that could be used as ‘good’ structures sheaves $\mathbf{A}$ in ADG. As Gel’fand Duality mandates, these are algebra sheaves over the suitably topologised spectrum of the original algebra.
    At the same time, at the back of Tasos’s mind, back in the day of the mid-90s when he was feverishly searching for solid founding pillars to erect ADG, must have surely been his beloved Differential Geometry [43]: that is to say, consciously or unconsciously (and here this author only speculates in retrospect), Tasos must have asked himself:
    $\bullet$ How can one extract differential geometric structure (not just topological) from Algebra (:algebraic structure), thus in a sense emulate Gel’fand duality, but in a differential geometric setting?
    To that end, motivating inspiration must have come to Tasos from another celebrated mathematical result, which also elegantly depicts the categorical duality between Algebra and Geometry: the Serre-Swan Theorem [76, 81], to which we briefly turn next.
  2. 2.
    The Serre-Swan Theorem: Jean-Pierre Serre’s version of the theorem, which to this author is more pertinent to Tasos’s original differential geometric endeavours and quests, roughly posits that for every commmutative unital (Noetherian) ring $R$, then the category of finitely generated projective $R$-modules (Algebra) is equivalent to the category of algebraic vector bundles $\mathcal{V}$ (i.e., locally free sheaves of structure sheaf $R$-modules of constant finite rank $n$) on the Spectrum $Spec(R)$ of $R$.
    The alert and astute reader, who is also familiar with the basic rudiments of ADG, must surely speculate that for Tasos, the result above must have come as an ‘epiphany moment’ in his quest for algebraic structures to model ADG, if one substitutes:
[…]
In toto, the usual $\mathcal{C}^{\infty}$-smooth vector bundles $\mathcal{V}$ on which the whole edifice of Classical Differential Geometry (CDG)—the so-called Classical Differential Calculus on Smooth (Differential) Manifolds—and its manifold applications to Modern Physics rest110110110We have empasised throughout our joint work with Tasos [53, 54, 55, 56] that CDG is a special case of, and can be recovered from, ADG when one assumes $\mathbf{A}\equiv\mathcal{C}^{\infty}(M)$—that is, when one simply assumes copies of the algebra $\mathcal{C}^{\infty}(M)$ of $\mathcal{C}^{\infty}$-smooth functions on a differential (spacetime) manifold $M$ as occupying the stalks of the structure sheaf $\mathbf{A}$ in the theory. Equivalently, it has been recently shown that the category $\mathcal{M}$ of $\mathcal{C}^{\infty}$-smooth manifolds is a full subcategory of the category $\mathcal{DT}$ of ADG-theoretic differential triads [9]. is abstracted and generalised in ADG by vector sheaves $\mathbf{\mathcal{E}}$ (i.e., locally free $\mathbf{A}$-modules of finite rank $n$) over an in principle arbitrary topological space $X$ [38, 39, 44].111111111Recall from Section 1 that in our finitary case, the base space on which our finitary sheaves of differential incidence Rota algebras $\Omega_{i}$ are soldered is the very (primitive) spectrum $Spec(\Omega_{i})$ of those algebras—a ‘discrete’ instance of ‘Gel’fand Duality meets Serre-Swan’.
As an additional bonus at this point, the ADG-theoretic extension and generalisation of the Serre-Swan Theorem was the principal move of Tasos in [40, 45, 44] towards arriving at a manifestly functorial geometric (pre)quantisation and second (:field) quantisation scheme for his ADG-theoretic field theory, with concomitant classification of the fields’ elementary particle quanta into bosons and fermions, as (32) above depicts.
3.1.2 Tracing the Physical ‘Origins’ of ADG: Breaking Algebra-Geometry Duality in Favour of an Algebraic Physical Dynamics
The two celebrated mathematical results mentioned above—Gel’fand Duality and the Serre-Swan Theorem—were seen to express a fundamental categorical duality (:functorial equivalence) between Algebra and Geometry and, arguably, they were speculated to be centers of inspiration and motivation for Professor Mallios in developing ADG as a mathematical abstraction, expansion, generalisation and enrichment112112112As Tasos originally preferred to call it cumulatively: an axiomatisation [38, 39]. of the usual CDG on smooth manifolds.
In this subsection, we turn our attention to Tasos’s fundamental Physical intuitions and motivations in applying ADG to what he used to call ‘Physical Space(Time)’ and ‘Physical Geometry’.
To that end, we quote directly from [48] the following ‘definition’ á-la Mallios of what he perceived as, and, in extenso, what ought to qualify as being, ‘Physical Geometry’:
…Physical Geometry is the ‘outcome’ of the physical laws…
For which he then displayed the following ‘causal nexus’ for producing Physical Geometry from Dynamical Laws:113113113Again, the quotation is borrowed verbatim from [43].
“…Now, by looking at the technical correspondence/association,
[…]
one realizes that [the displayed expression above] might also be construed, as an equivalent analogue of the implication;
[…]
Consequently, still to repeat the above, but state it otherwise, one concludes that:
It is actually the physical laws, that make, what we might call (physical) ‘geometry’….”
All of Mallios’s prophetic musings above may be subsumed under the following distilled Fundamental Aphorism:
Fundamental Aphorism: Physical Space(Time)’ and ‘Physical Geometry’ is the result or the product of Field Dynamics, in much the same way that, as we saw earlier in the paper, the Curvature Field is a Geometric Morphism image of the Connection Field.
In this line of thought, we conclude the present paper by quoting verbatim below the very opening paragraph of the wonderful Exordium in Mallios and Zafiris’s last joint research monograph [61]:114114114Some emphasised parts in the quote above are our emphasis.
The major aim of the application of Analysis and Differential Geometry in Physics is the setting up of a mechanism providing a precise description of the emergence of geometric spectrums115115115That is to say, ‘geometric spaces’. due to dynamical interactions, which can be further used for making predictions. In this manner, the notion of a geometric spectrum is considered as the outcome of a physical law or more generally of a dynamical interaction of a particular form. This raises immediately the question if there exists an approach to physical geometric spectrums that is independent of any coordinate point manifold background, in the sense that it refers directly to the physical relations causing the appearance of these spectrums without the intervention of any ad hoc coordinate choices. The answer provided to this question in this book is that the theory of differential vector sheaves, that is geometric vector sheaves equipped with a connectivity structure and obeying appropriate cohomological conditions, provides the sought after functorial tool for a universal and natural approach to physical geometric spectrums. The major difference of the proposed approach in comparison to the traditional ones based on classical differential calculus and differential geometry of smooth manifolds consists in the realization that a classical analytic technique is susceptible to a natural background-independent generalization if it is localizable by sheaf-theoretic means. In this case the technique can be expressed functorially, that is by means of natural transformations of sheaf functors via the machinery of homological algebra. This is of crucial significance for setting up a mechanism describing the emergence of physical geometric spectrums where the notion of background smoothness is inapplicable…”
3.2 A Wish for the Future
The whole time that I have known him, Tasos was always with the underdog, “always supporting and taking sides with the hunted, not the hunter”,116116116One of his own ‘proverbs of wisdom’: “We are always with the hunted, not the hunter”. which he kept reminding me on a day-to-day basis. He also constantly urged me to take risks and be unconventional, unorthodox and iconoclastic in my research quests and endeavours in Quantum Gravity [71].
Thus, in paying tribute and homage to my wonderful teacher, mentor, friend and immortal companion in our joint Unending Quest, here’s one of Tasos’s favourite Feynman Quotes:117117117Taken from Richard Feynman’s intro to his [15], where he talks about researchers taking the risk and ‘going off into the wild blue yonder’ realm, seemingly strange and largely unexplored yet landscape, of Quantum Gravity.
“…It is important that we don’t all follow the same fashion. We must increase the amount of variety and the only way to do this is to implore you few guys, to take a risk with your own lives so that you will never be heard of again, and go off into the wild blue yonder to see if you can figure it out…”
May his far reaching vision, the breadth of his conceptual perception, the imagination of his mathematical and physical intuition, the depth of his philosophical enquiry, the originality and the unorthodoxy of his approach, as well as the risk and the adventurousness of his research endeavours—all coupled to the priceless legacy that Professor Anastasios Mallios leaves behind—nurture, enrich, motivate and inspire future researchers in Quantum Gravity for years to come!
Addendum: An Anecdotal Exchange with Professor Mallios and a Conclusion Drawn from It
I would like to share with this forum a private two-part exchange that we enjoyed with Tasos way back in May 1998, actually on the day of my 30th birthday, more than two years before his first 2-volume pitch of Abstract Differential Geometry: The Geometry of Vector Sheaves was published by Kluwer Academic Publishers [38]. That exchange on the one hand dovetails perfectly with the two main mathematical results mentioned in the last section118118118Gel’fand Duality and the Serre-Swan theorem. that, as I argued, must have played a pivotal role, consciously and/or unconsciously, in motivating and inspiring Tasos to develop ADG in the first place, and on the other, it casts light on Tasos’s character as a pure and young-at-heart, decent human being.
I had just obtained my Ph.D. [66], with Professor Mallios as one of the external examiners of my thesis, and, very interested in the early developments of Topos Theory (TT) [37], I was naturally attracted by the core mathematical ideas and the working philosophy of one of TT’s main early architects: Alexandre Grothendieck [24]. I had also just finished reading the first part (:Fatuity and Renewal) of Grothendieck’s ‘autobiographical’ manuscript titled Reaping and Sowing [25] and we were discussing the wide range and far reaching depth of Alexandre Grothendieck’s contributions to Modern Mathematics, especially to the field of Algebraic Geometry, via the introduction and application of novel Homological Algebra (:Category-theoretic) ideas, concepts and technical constructions.
Part 1 of the Exchange: Grothendieck’s ‘Working Philosophy’ I remember I initiated the exchange by telling Tasos that I had just finished reading the first part of the Récoltes and made the remark that the gist of Grothendieck’s working philosophy in Mathematics was to attain an epoptic—as broad, as general and as abstract—viewpoint of the entire landscape of Mathematics. Tasos agreed and added two crucial ingredients:
  1. 1.
    That Grothendieck, willingly or not, explicitly or not, formally or informally (:intuitively), essentially axiomatised Mathematics; and,
  2. 2.
    That Grothendieck used to encounter and stare at the (only apparently) complex and esoteric Mathematics’ landscape with the innocence and the ignorance of a child.
$\bullet$ He backed the first point by saying that, much in the same vain as Grothendieck, he was planning to call (as he actually did!) his forthcoming work on ‘Abstract Differential Geometry: The Geometry of Vector Sheaves’ an Axiomatic Approach to Differential Geometry [38]. I asked him why did he think that the most epoptic, bird’s eyeview of Mathematics could be attained by Axiomatisation, and he quoted me Aristotle from Nicomachean Ethics: “He who can properly define, divide and distinguish is to be considered God[1]. He said that in Mathematics one can attain the broadest and most-encompassing viewpoint only when one has properly clarified and laid down the fundamental concepts and base axioms.
$\bullet$ He backed his second point by quoting back to me Bertrand Russell:119119119This is one of Tasos’s Top-3 quotes.Men are born ignorant, not stupid. Education makes them stupid.”. He said that Grothendieck had the gift and ability to look at the World of Mathematics in innocent and ignorant awe and amazement, with fresh eyes, not at all conditioned or biased by prior Knowledge or Education, but perhaps more importantly, by not being afraid of making mistakes.
I indeed recall reading, a couple of days earlier, from the first part of Grothendieck’s Récoltes the following telling excerpt:
…Discovery is the privilege of the child. It’s the little child that I want to talk about, the child who is not afraid to be wrong, to look silly, to not be serious, to not be like everyone else. He is neither afraid that the things he looks at will have a bad taste, different from what he expects, from what they appear to be, or rather: from what he has already understood them to be. He ignores the unspoken and unwavering consensus that form part of the air we breathe - which all the grown-ups are supposed to know and they do know. God knows (I suppose the grown-ups know well) if there have been any such child, since the dawn of ages!…
…The little child discovers the world as he breathes - the ebb and flow of his breath make him welcome the world in its delicate being, and makes him project himself into the world that also welcomes him. The adult can also discover, in those rare moments when he has forgotten his fears and his knowledge, when he looks at things or himself with eyes wide open, eager to know, new eyes - the eyes of a child…
Part 2 of the Exchange: Grothendieck’s Work With regard to Grothendieck’s main contribution to Mathematics, especially in Algebraic Geometry, Tasos maintained that Grothendieck essentially abstracted and purely algebraicised Algebraic Geometry, and effectively substituted Hard Analysis on ‘Rigid’ Geometrical (:Arithmetic) Spaces by the more malleable and flexible inherently algebraic concepts and methods of Sheaf Cohomology.
With the advent, blossoming and effective manifold applications of ADG to fundamental Mathematics and Physics, I can now draw confidently the following parallel in posthumous honour of Tasos:
What Grothendieck did for Algebraic Geometry, Mallios did for Differential Geometry.
Acknowledgments
After a hiatus of one-and-a-half decades, this paper comes as a result of serendipitous privilege and dogged perseverance: serendipitous privilege for crossing worldlines and working closely with Tasos Mallios, and dogged perseverance from having read, on the year that Tasos departed, Charles Bukowski’s poem Go All the Way. Both gentlemen taught me in their own inimitable way that “Isolation is the Gift[8], for which I am eternally indebted to them.
The unceasing ‘moral’ support of my lovely family: Kathleen, Francis, James and Cookie, is also lovingly aknowledged, especially their patience and understanding in putting up with me over the years.